### v1

# Toward Relative Positional Encoding in Spiking Transformers

Abstract
Spiking neural networks (SNNs) are bio-inspired networks that model how neurons in the brain communicate through discrete spikes, which have great potential in various tasks due to their energy efficiency and temporal processing capabilities. SNNs with self-attention mechanisms (Spiking Transformers) have recently shown great advancements in various tasks such as sequential modeling and image classifications. However, integrating positional information, which is essential for capturing sequential relationships in data, remains a challenge in Spiking Transformers. In this paper, we introduce an approximate method for relative positional encoding (RPE) in Spiking Transformers, leveraging Gray Code as the foundation for our approach. We provide comprehensive proof of the method’s effectiveness in partially capturing relative positional information for sequential tasks. Additionally, we extend our RPE approach by adapting it to a two-dimensional form suitable for image patch processing. We evaluate the proposed RPE methods on several tasks, including time series forecasting, text classification, and patch-based image classification. Our experimental results demonstrate that the incorporation of RPE significantly enhances performance by effectively capturing relative positional information.
Spiking Neural Networks, Relative Positional Encoding
1 Introduction
Spiking Neural Networks (SNNs) (Maass, 1997) are a class of bio-inspired models designed to emulate the communication process of biological neurons, which transmit information through discrete spikes. In contrast to artificial neural networks (ANNs) that operate on continuous values, SNNs process information in the form of spikes occurring at precise moments in time. This discrete event-driven processing enables SNNs to be highly energy-efficient. The temporal characteristics of spikes make SNNs particularly well-suited for tasks involving sequential data or dynamic environments, such as sensory processing (Li et al., 2017a; Fang et al., 2021), patch-based image classification (Zhou et al., 2023b; Yao et al., 2023), time-series forecasting (Lv et al., 2024b; Park et al., 2024), and natural language processing (Zhu et al., 2023; Lv et al., 2023b; Xing et al., 2024).
In ANNs, relative positional encoding (RPE) is often used to encode the positions of input elements (Su et al., 2024), enabling the model to capture the relative relationships between positions within a sequence. While RPE has been shown to enhance performance in language modeling (Su et al., 2024) and image processing (Liu et al., 2021), its direct application to SNNs remains limited due to the unique characteristics of spike-based computation. Current approaches to realize positional encoding (PE) in SNNs either fail to maintain the uniqueness of spike representations for each position (Zhou et al., 2023b; Yao et al., 2024a), or ignore the incorporation of relative positional information (Lv et al., 2024a).
In this paper, we first identify a limitation in the current spiking self-attention (SSA) mechanism (Zhou et al., 2023b), which is suboptimal for measuring the distance (or similarity) between tokens in the query and key matrices. Specifically, the dot product operation fails to effectively differentiate distances when both the query and key are spike matrices. To address this limitation, we propose a new spiking self-attention mechanism based on the Exclusive-NOR (XNOR) operation, which enhances the measurement of query-key similarity by leveraging the Hamming distance (see Section 3.4).
Secondly, we introduce a novel approximate relative positional encoding method for SNNs, based on Gray Code (Gray, 1953) as the foundational encoding scheme. The consistent Hamming distance between Gray Code representations of two position indices differing by $2^{n}$ allows for serving as a spiking RPE method and the integration with the XNOR-based SSA mechanism. The binary encoding nature of Gray Code is well-suited to represent relative positions while preserving the spiking dynamics of SNNs, which is crucial for effectively processing sequential data. Furthermore, we provide a theoretical proof in Appendix A demonstrating the efficacy of Gray Code in partially capturing relative positional information for sequential tasks.
Moreover, we extend our RPE method to the logarithmic form that can be directly integrated into the attention map of XNOR-based spiking self-attention, offering greater efficiency compared to the previous form while maintaining comparable performance. For image patches, we also design a two-dimensional version of our method to capture the relative positional information of both height and width.
To evaluate the effectiveness of our RPE approaches, we apply them to three diverse tasks: time series forecasting, text classification, and patch-based image classification. Our experiments demonstrate that the proposed RPE method enhances the performance of SNNs across all three tasks, allowing the original SNNs to capture and leverage relative relationships between elements in the input sequential data. Moreover, we further conduct an ablation study and positional encoding analysis to validate the inner properties of our proposed RPE method. Based on the analysis, we offer several valuable insights into designing effective positional encoding schemes for SNNs.
We believe that this work contributes a new method for integrating relative positional encoding into SNNs, expanding their potential for a wide range of applications in machine learning and computational neuroscience. To summarize, our contributions are as follows:
  • A Novel RPE Method for SNNs. Our research represents one of the pioneering efforts to propose a solution to relative positional encoding for spiking neural networks, albeit as an approximation method.
  • Consistent Performance Improvement. Our proposed method and its extended variants significantly and consistently improve the performance of SNNs across various sequential tasks, including time-series forecasting and text classification.
  • Theoretical Validation and Insightful Analysis. In addition to experimental evaluations, we provide thorough theoretical proof to validate the effectiveness of our RPE method. Furthermore, we offer an analysis of the underlying properties of RPE, providing valuable insights for future designs of RPE in SNNs.
2 Related Work
In sequential tasks, positional encoding (PE) plays a crucial role in capturing the order of tokens. Traditional absolute positional encoding assigns a fixed position to each token, whereas relative positional encoding (RPE) emphasizes the distance between tokens. RPE allows the model to generalize across different sequence lengths and better capture relationships between tokens.
Despite its significance, there has been limited research focusing on the role of positional encoding approaches in spiking neural networks. Notably, both Spikformer (Zhou et al., 2023b) and Spike-driven Transformer (Yao et al., 2023, 2024a) incorporate a combination of convolutional layers, batch normalization, and spiking neuron layers to generate “learnable relative positional encodings”. However, we argue that this approach functions more similarly to a spike-element-wise residual connection (Fang et al., 2021) than to a conventional positional encoding module. A robust PE module should offer unique representations for each position, but the spike-position matrices generated by these methods may lead to identical spike representations for different positions.
Lv et al. (2024a) introduced a spiking positional encoding method for SNNs, inspired by central pattern generators (Marder & Bucher, 2001). This method generates periodic positional encoding, ensuring a unique binary spike matrix for each position. However, their approach is based on absolute positional encoding and, thus, does not capture the time-translational invariance property in many sequential modeling problems, which, however, is an important advantage of relative positional encoding methods.
3 Preliminary
Figure 1: Preliminary Figures. (a) Spike dynamics of spiking (LIF) neurons. (b) Illustration of spiking self-attention in Spikformer (Zhou et al., 2023b) and QKFormer (Zhou et al., 2024). (c) An example of Hamming Distance. (d) Calculation process of Gray Code.
3.1 Spiking Neural Networks
The fundamental building block of SNNs is the spiking neuron, such as the leaky integrate-and-fire (LIF) model (Maass, 1997), which is governed by the input current $I(t)$, influencing the membrane potential $U(t)$ and the spike output $S(t)$ at each time step $t$. The dynamics of the LIF neuron are captured by the following system of equations:
$U(t)=H(t-\Delta t)+I(t),\quad I(t)=f(\mathbf{x};\mathbf{\theta}),$ (1)
$H(t)=V_{\rm reset}S(t)+(1-S(t))\beta U(t),$ (2)
$S(t)=\begin{cases}1,&\text{if }U(t)\geq U_{\rm thr},\\ 0,&\text{if }U(t)<U_{\rm thr}.\end{cases}$ (3)
Here, $I(t)$ represents the spatial input to the LIF neuron at time $t$, which is computed by the function $f$ using the input vector $\mathbf{x}$ and the learnable parameters $\mathbf{\theta}$. The discretization constant $\Delta t$ controls the time granularity in the LIF model, and $H(t)$ is the temporal output at time step $t$. The spike $S(t)$ is determined by a Heaviside step function that is activated when the membrane potential $U(t)$ exceeds a threshold $U_{\rm thr}$. Upon reaching the threshold, the neuron generates a spike, and $H(t)$ resets to $V_{\rm reset}$. If $U(t)$ does not exceed the threshold, the membrane potential decays to $H(t)$ at a rate determined by $\beta$.
3.2 Spiking Self-Attention
Spiking self-attention (SSA) is a spiking version of self-attention (Vaswani, 2017), which was proposed in Spikformer (Zhou et al., 2023b) and QKFormer (Zhou et al., 2024). The vital design is to utilize discrete spikes to approximate the vanilla self-attention mechanism. It can be written as:
$\small\begin{split}\mathbf{Q}&=\mathcal{SN}\left(\operatorname{BN}\left(\mathbf{X}\cdot\mathbf{W_{Q}}\right)\right),\\ \mathbf{K}&=\mathcal{SN}\left(\operatorname{BN}\left(\mathbf{X}\cdot\mathbf{W_{K}}\right)\right),\\ \mathbf{V}&=\mathcal{SN}\left(\operatorname{BN}\left(\mathbf{X}\cdot\mathbf{W_{V}}\right)\right),\end{split}$ (4)
where $\mathcal{SN}$ is a spike neuron layer described in Equation 3. The input is denoted as $\mathbf{X}\in\{0,1\}^{T\times L\times D}$, where $T$ is the number of time steps. $\operatorname{BN}$ represents batch normalization, and $\tau$ is a scaling factor. The attention map $\mathbf{AttnMap}$ is then computed as the dot product between $\mathbf{Q}$ and $\mathbf{K}^{\operatorname{T}}$:
$\small\mathbf{AttnMap}=\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}},$ (5)
$\small\operatorname{SSA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)=\mathcal{SN}(\operatorname{BN}((\mathbf{Attn}\cdot\mathbf{V}*\tau)\cdot\mathbf{W})).$ (6)
As a result, the attention map $\mathbf{AttnMap}\in\mathbb{N}_{0}^{T\times L\times L}$, where $\mathbb{N}_{0}$ denotes the set of non-negative integers. The outputs of the SSA, as well as $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, are all spike matrices containing only values of 00 and $1$. The parameters $\mathbf{W_{Q}}$, $\mathbf{W_{K}}$, $\mathbf{W_{V}}$, and $\mathbf{W}$ are all learnable parameters.
3.3 Relative Positional Encoding
In self-attention mechanisms, the attention weights are modified using a relative positional encoding (RPE) $\mathbf{R}_{i,j}$, which depends on the distance between the $i$-th and $j$-th tokens. A typical realization of RPE (Liu et al., 2021; Hao et al., 2024) is formulated as:
$\small\begin{split}\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})&=\mathbf{AttnMap\cdot V}\\ &=\text{Softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}}}{\sqrt{d_{k}}}+\mathbf{R}_{i,j}\right)\mathbf{V}.\end{split}$ (7)
Here, $\mathbf{R}_{i,j}$ represents the relative distance between tokens $i$ and $j$. Specifically, the relative positional encoding for the $i$-th position in $\mathbf{Q}$ and the $j$-th position in $\mathbf{K}$ should satisfy distance consistency, i.e., the magnitude of $\mathbf{R}_{i,j}$ should depend solely on the relative distance $|i-j|$, enabling the model to distinguish between nearby and distant tokens. This approach improves the model’s ability to handle long-range dependencies and sequence variations.
Figure 2: Overview of Our Method. (a) XNOR-based Spiking Self-Attention (SSA). This panel illustrates the computation flow for the query ($\mathbf{Q}$) and key ($\mathbf{K}$) matrices in a PyTorch-style notation. (b) Gray-PE. Gray code representations exhibit a consistent Hamming distance for two integers differing by $2^{n}$. Gray-PE is implemented by concatenating $G(\bm{l})$ along the $D$ dimension of both $\mathbf{Q}$ and $\mathbf{K}$. (c) Log-PE. A pre-assigned relative distance encoding map $\mathbf{R}_{i,j}\in\mathbb{N}_{0}$ is added to the standard attention map $\mathbf{AttnMap}$. The shading indicates the magnitude of the values, with deeper blue corresponding to larger values. (d) 2D Form of Gray-PE. For image patches, a 2D relative positional encoding is more suitable than the 1D version, as it better captures the spatial relationships.
3.4 Hamming Distance
The Hamming distance between two binary strings of equal length is the number of bit positions at which the corresponding bits are different. Formally, for two binary strings $A$ and $B$ of length $m$,
$\small d_{H}(A,B)=\sum_{i=1}^{m}\delta(A_{i},B_{i}),$ (8)
where
$\small\delta(A_{i},B_{i})=\begin{cases}1&\text{if }A_{i}\neq B_{i},\\ 0&\text{otherwise}.\end{cases}$ (9)
Since the preliminary knowledge involved is extensive and loosely connected, we have provided Figure 1 to help readers visually grasp the key concepts of each section.
4 Method
4.1 Gray Code Definition
Gray Codes (Gray, 1953), also known as reflected binary codes, are binary sequences where consecutive numbers differ by exactly one bit. Gray Code $G(x)$ for a non-negative integer $x$ can be generated using the formula: The standard binary reflected Gray Code $G(x)$ for a non-negative integer $x$ is defined as:
$\small G(x)=x\oplus\left(x\gg 1\right),$ (10)
where $\oplus$ denotes the bitwise XOR operation, and $\gg$ denotes the arithmetic right shift (with zero padding).
4.2 XNOR-Based Spiking Self-Attention
In the standard self-attention mechanism of Transformer (Vaswani, 2017), the attention map is computed as the dot product between the query matrix $\mathbf{Q}$ and key matrix $\mathbf{K}$, i.e., $\mathbf{AttnMap}=\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}}$. The primary concept behind this approach is to use the dot product to measure the similarity between each token in $\mathbf{Q}$ and $\mathbf{K}$, which works effectively when both $\mathbf{Q}$ and $\mathbf{K}$ are continuous-valued matrices. However, when $\mathbf{Q}$ and $\mathbf{K}$ are spike matrices, the dot product spike matrices, the dot product becomes inadequate for capturing the temporal or event-based interactions intrinsic to spike-based representations. As a result, the dot product operation in spiking self-attention, as outlined in Section 3.2, is not well-suited for spiking Transformers. To address this limitation, several researchers (Yao et al., 2023; Zhou et al., 2023a; Yao et al., 2024b) have proposed alternative attention mechanisms specifically designed for SNNs.
In this paper, we propose an XNOR-based spiking self-attention mechanism, which leverages the Hamming distance (See Section 3.4) to measure the similarity between each token in $\mathbf{Q}$ and $\mathbf{K}$. This approach enables the integration of the Gray-code-based relative positional encoding method into the attention mechanism, allowing for more effective capture of positional information in spike-based representations. Formally, we modify Equation 5 as follows:
$\small\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot(\mathbf{Q}\oplus\mathbf{K}),$ (11)
where $\lnot$ denotes the Not operation, $\oplus$ denotes the XOR operation, and $D$ represents the channel dimension. It is important to note that every token in $\mathbf{Q}$ will perform XOR $\oplus$ with every token in $\mathbf{K}$, so we sum over the channel dimension $D$ to get $\mathbf{AttnMap}\in\mathbb{N}_{0}^{T\times L\times L}$, shown in Figure 2 (a). Compared to the dot product, the XNOR operation significantly increases the values in $\mathbf{AttnMap}\in\mathbb{N}_{0}$. Therefore, the scale factor $\tau$ in Equation 6 should be set to a smaller value or treated as a learnable parameter, ensuring that the firing rate of $\mathcal{SN}$ does not become excessively large. We will empirically demonstrate in the Experiments Section that this XNOR modification does not negatively impact the performance of the vanilla spiking self-attention.
4.3 Approximate RPE Method Based on Gray Code
We modify the SSA module by replacing the dot product operation with the XNOR operation, enabling the use of Hamming distance to measure the similarity between tokens. In this context, we propose that Gray Code can serve as an approximate approach to relative positional encoding for SNNs. This is supported by Theorem 4.1, which demonstrates the effectiveness of Gray Code in partially capturing positional relationships within the framework of XNOR-based spiking self-attention.
Theorem 4.1.
For two integers differing by $2^{n}$ in decimal, their Gray Code representations have a consistent Hamming distance of $1$ for $n=0$, and $2$ for $n\geq 1$, i.e., for any non-negative integer $a$ and $n$:
$d_{H}(G(a),G(a+2^{n}))=\begin{cases}1&\text{if }n=0,\\ 2&\text{if }n\geq 1.\end{cases}$ (12)
Proof.
Please refer to Appendix A for detailed proof. ∎
From Theorem 4.1, we can infer that Gray Code ensures the consistency of distance representations for every $2^{n}$ ($n\geq 0$) relative distance. This property makes Gray Code an ideal approximate RPE method for SNNs, as it not only partially captures relative positions but also preserves the binary nature (00 and $1$) nature of SNNs.
In our implementation, we concatenate the Gray Code representations of each position index to both the query matrix $\mathbf{Q}\in\mathbb{N}_{0}^{T\times L\times D}$ and key matrix $\mathbf{K}\in\mathbb{N}_{0}^{T\times L\times D}$, leaving the remaining operations unchanged. Formally, the attention map $\mathbf{AttnMap}$ is computed as follows:
$\small\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot([\mathbf{Q}\parallel G(\bm{l})]\oplus[\mathbf{K}\parallel G(\bm{l})]),$ (13)
where $G(\cdot)$ represents the function that converts non-negative integers into their corresponding Gray Code representations. The vector $\bm{l}$ denotes an array of position indexes, specifically $[0,1,2,\dots,L-1]$, where $L$ is the sequence length of both $\mathbf{Q}$ and $\mathbf{K}$. $\parallel$ denotes concatenation on the channel dimension $D$. We name this method as Gray-PE.
4.4 Extended Variant on Attention Maps
Although Gray-PE can partially capture relative distances, it faces significant challenges when the input sequence length is large or when the downstream task is highly sensitive to long-range dependencies. For instance, when the input sequence length reaches $10^{2}$, Gray-PE can only distinguish a limited range of relative distances. To address this issue, we propose an extension of Gray-PE, referred to as Log-PE. Specifically, we simulate Equation 7 by directly adding a pre-assigned relative position map, denoted as $\mathbf{R}_{ij}$, to the attention map produced by the XNOR-based SSA operation. To be formal:
$\small\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot(\mathbf{Q}\oplus\mathbf{K})+\mathbf{R}_{i,j}$ (14)
where
$\mathbf{R}_{i,j}=\begin{bmatrix}R_{i,j}\end{bmatrix}=\begin{bmatrix}\lceil\log_{2}(\frac{L-1}{|i-j|})\rceil\end{bmatrix}.$ (15)
Here, $\lceil$.$\rceil$ . ⌉ denotes the round-up function, $L$ is the sequence length, and $i,j$ Figure 2 (c) shows an illustration of Log-PE, indicating that the farther relative distance, the smaller $R_{i,j}$. Since the original $\mathbf{AttnMap}\in\mathbb{N}_{0}$ is a matrix composed of non-negative integers, our primary goal is to ensure accurate relative distance consistency while preserving the effectiveness of spiking self-attention. If we set the $R_{i,j}$ as $\frac{L-1}{|i-j|}$, we could obtain a complete relative positional encoding for the XNOR-based SSA. However, we choose not to pursue this solution because, for long sequence lengths $L$, the large values of $\frac{L-1}{|i-j|}$ would severely distort the original representations in the spiking self-attention map. Therefore, using the logarithmic form $R_{i,j}$ represents a trade-off that balances the values of the attention map and relative positional encoding.
4.5 Two-Dimensional Form for Image Patches
CNN-based SNN models, such as Spiking VGG (Sengupta et al., 2019) and SEW-ResNet (Fang et al., 2021), do not incorporate the concept of “positional encoding” in their spike representations. In contrast, Vision Transformer (Dosovitskiy et al., 2021) reformulated traditional image classification into a patch-based approach, dividing images into smaller patches. Unlike 1D positional encoding, which only considers the linear sequence of patches, 2D RPE accounts for both the horizontal and vertical positions of the patches in the image grid. This ensures that the model can recognize the relative positions along a single axis and the crucial interactions between patches across both dimensions. In our implementation, we assign horizontal and vertical positions with independent dimensions to store the Gray Code. Formally, the attention map $\mathbf{AttnMap}$ is:
$\small\begin{split}\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot(&[\mathbf{Q}\parallel G(\mathbf{h})\parallel G(\mathbf{w})]\\ \oplus&[\mathbf{K}\parallel G(\mathbf{h})\parallel G(\mathbf{w})]).\end{split}$ (16)
Here, $\mathbf{h}$ is the array of position indexes, specifically $\mathbf{h}=[0,1,2,\dots,h-1]$, where $h$ denotes the maximum patch index along the height axis. Similarly, $\mathbf{w}$ is along the width axis. We show our 2D form in Figure 2 (d).
Table 1: Experimental results of time-series forecasting on $4$ benchmarks with various prediction lengths $6,24,48,96$. “PE” stands for positional encoding. “R” denotes relative PE, while “A” denotes absolute PE. “w/o” denotes “without” while “w/” denotes “with”. The best results for each series of SNNs are highlighted in bold font. $\uparrow$ (indicating "higher is better") and $\downarrow$ (indicating "lower is better") are used to denote performance metrics. Results highlighted with shading are ours. All results are averaged across $3$ random seeds.
Models PE Metric Metr-la (Short) Pems-bay (Short) Solar (Long) Electricity (Long) Avg.
Spike Type $6$ $24$ $48$ $96$ $6$ $24$ $48$ $96$ $6$ $24$ $48$ $96$ $6$ $24$ $48$ $96$
Transformer w/ RPE R R2$\uparrow$ .$\bf.729$$\bf.560$$\bf.416$$\bf.306$$\bf.787$$730$$\bf.694$$\bf.676$$951$$854$$\bf.763$$\bf.720$$\bf.984$$\bf.978$$\bf.974$$\bf.968$$\bf.756$
[…]
Transformer w/ Sin-PE A R2$\uparrow$ .$727$$554$$413$$284$$785$$\bf.734$$688$$673$$\bf.953$$\bf.858$$759$$718$$978$$975$$972$$964$752.752.752.752
[…]
Spikformer w/o PE R2$\uparrow$ .$697$$491$$383$$242$$768$$684$$678$$663$$903$$819$$715$$656$$956$$955$$953$$943$719.719.719.719
[…]
Spikformer w/ Conv-PE (Original) A R2$\uparrow$ .$713$$527$$399$$267$$773$$697$$686$$667$$929$$828$$744$$674$$959$955.955.955.955 .955.955.955.955 .$954$733.733.733.733
[…]
Spikformer w/ CPG-PE R2$\uparrow$ .$726$$526$$419$$287$$780$$712$$690$$666$$937$$833$$757$$707$$972$$970$$966$$960$744.744.744.744
[…]
Spikformer-XNOR w/o PE R2$\uparrow$ .$699$$488$$380$$244$$764$$680$$679$$666$$906$$820$$721$$653$$960$$958$$951$$943$720.720.720.720
[…]
Spikformer-XNOR w/ Conv-PE A R2$\uparrow$ .$718$$531$$405$$269$$771$$693$$690$$665$$928$$829$$740$$669$$960$$957$$955$$953$733.733.733.733
[…]
Spikformer-XNOR w/ Binary-PE R2$\uparrow$ .$701$$488$$384$$252$$759$$698$$675$$644$$918$$815$$707$$677$$964$$959$$949$$945$721.721.721.721
[…]
R2$\uparrow$ .$728$$\bf.544$$414$$\bf.295$$782$$\bf.724$$\bf.694$$\bf.673$$\bf.936$$840$$756$$710$$974$$972$$966$$962$748.748.748.748
[…]
R2$\uparrow$ .$\bf.735$$535$$\bf.424$$290$$\bf.789$$717$$691$$670$$933$$\bf.841$$\bf.758$$\bf.734$$\bf.978$$\bf.974$$\bf.968$$\bf.964$$\bf.750$
[…]
QKFormer w/o PE (Original) R2$\uparrow$ .$702$$501$$377$$240$$750$$702$$664$$655$$905$$751$$497$$375$$964$$963$$958$$955$685.685.685.685
[…]
QKFormer w/ Conv-PE A R2$\uparrow$ .$717$$513$$376$$246$$767$$706$$681$$654$$920$$748$$512$$416$$970$$967$$963$$958$695.695.695.695
[…]
QKFormer w/ CPG-PE A R2$\uparrow$ .$740$$554$$419$$276$$783$$714$$702$$660$$922$$754$$702$$604$$977$$969$$968$$963$732.732.732.732
[…]
R2$\uparrow$ .$\bf.742$$\bf.551$$\bf.418$$\bf.274$$799$$\bf.715$$691$$\bf.674$$927$$817$$710$$691$$974$$970$$968$$965$742.742.742.742
[…]
R2$\uparrow$ .$\bf.742$$541$$416$$265$$\bf.801$$710$$\bf.707$$661$$\bf.928$$\bf.818$$\bf.748$$\bf.698$$\bf.978$$\bf.974$$\bf.972$$\bf.966$$\bf.746$
[…]
5 Experiments
5.1 Datasets
To evaluate the RPE capabilities of the compared models, we conduct experiments on two sequential tasks: time-series forecasting and text classification. Following Lv et al. (2024b), we choose $4$ real-world datasets for time-series forecasting: Metr-la (Li et al., 2017b), Pems-bay (Li et al., 2017b), Electricity (Lai et al., 2018), Solar (Lai et al., 2018). For text classification, we follow Lv et al. (2024a) and conduct experiments on six benchmark datasets: Movie Reviews (Pang & Lee, 2005), SST-2 (Socher et al., 2013), SST-5, Subj, ChnSenti, and Waimai. Additionally, to demonstrate the versatility of our RPE method in image processing, we perform patch-based image classification experiments on one static dataset, CIFAR, and one neuromorphic dataset, CIFAR10-DVS (Li et al., 2017a). The details of these datasets, metrics, and training hyper-parameters are provided in Appendix B.
5.2 Time-Series Forecasting
Time-series forecasting is a suitable task for evaluating the effectiveness of our proposed RPE method because it involves capturing the temporal dependencies between data points over time. We follow the framework outlined by Lv et al. (2024b) to conduct time-series forecasting experiments using SNN models. Specifically, we use Spikformer (Zhou et al., 2023b) and the current state-of-the-art (SOTA) model, QKFormer (Zhou et al., 2024), as the backbone architectures, both of which incorporate spiking self-attention mechanisms. We modify the spiking self-attention mechanism as outlined in Section 4.2 to create two variants: Spikformer-XNOR and QKFormer-XNOR. Subsequently, we integrate our proposed Gray-PE and its extended forms into these SNN architectures.
For the baseline comparisons, we consider the following: (1) spiking absolute PE and CPG-PE (Lv et al., 2024a), (2) the PE used in the original architectures, referred to as Conv-PE, and (3) the integration of Binary-PE into the XNOR-based SSA. To isolate the effects of the XNOR modification, we also perform experiments on Spikformer-XNOR and QKFormer-XNOR without any positional encoding, as well as with Conv-PE. We present the performance of the compared SNN models with various positional encoding methods in Table 1. The key findings are as follows:
Table 2: Accuracy (%) on $6$ text classification benchmarks. The best results of SNNs and ANNs are formatted in bold font. We set the sentence length to $128$ for the English dataset and $32$ for the Chinese dataset. Experimental results are averaged across $5$ random seeds.
Model PE Param (M) English Dataset (Length = $128$) Chinese Dataset (Length = $32$) Avg.
Spike Type MR SST-2 Subj SST-5 ChnSenti Waimai
Fine-tuned BERT A $109.8$ $\bm{87.63}$±$\pm 0.18$ $\bm{92.31}$±$\pm 0.17$ $\bm{95.90}$±$\pm 0.16$ $\bm{50.41}$±$\pm 0.13$ $\bm{89.48}$±$\pm 0.16$ $\bm{90.27}$±$\pm 0.13$ $\bm{84.33}$
Spikformer w/o PE $109.8$ $75.87$±$\pm 0.35$ $81.71$±$\pm 0.31$ $91.60$±$\pm 0.30$ $41.84$±$\pm 0.39$ $85.62$±$\pm 0.25$ $86.87$±$\pm 0.28$ $77.25$
Spikformer w/ CPG-PE A $110.4$ $82.42$±$\pm 0.42$ $82.90$±$\pm 0.33$ $92.50$±$\pm 0.25$ $43.62$±$\pm 0.36$ $86.54$±$\pm 0.26$ $\bf 88.49$±$\pm 0.29$ $79.41$
Spikformer-XNOR w/o PE $109.8$ $75.80$±$\pm 0.40$ $81.74$±$\pm 0.40$ $91.50$±$\pm 0.29$ $41.88$±$\pm 0.38$ $85.64$±$\pm 0.31$ $86.66$±$\pm 0.33$ $77.20$
Spikformer-XNOR w/ Gray-PE R $109.8$ $83.73$±$\pm 0.45$ $84.52$±$\pm 0.39$ $92.50$±$\pm 0.33$ $44.06$±$\pm 0.48$ $87.41$±$\pm 0.36$ $88.40$±$\pm 0.30$ $80.11$
Spikformer-XNOR w/ Log-PE R $109.8$ $\bf 83.88$±$\pm 0.40$ $\bf 84.64$±$\pm 0.37$ $\bf 92.80$±$\pm 0.30$ $\bf 44.52$±$\pm 0.43$ $\bf 87.95$±$\pm 0.34$ $88.46$±$\pm 0.28$ $\bf 80.38$
(1) Relative PE generally outperforms absolute PE in time-series forecasting tasks. Based on the comparison between the Transformer with RPE and sinusoidal-PE, as well as Spikformer-XNOR with Conv-PE and Gray-PE, we observe that relative PE consistently outperforms absolute PE, particularly on long-term time-series datasets such as Solar and Electricity.
(2) The XNOR modification does not impact the performance of the original SNN models. The average performance of Spikformer without PE is nearly identical to that of Spikformer-XNOR without PE. This suggests that our XNOR modification of the SSA does not affect the performance of the original SNN models. Consequently, we can attribute the observed performance improvements solely to the introduced positional encoding techniques, excluding any impact from the XNOR operation itself.
(3) Gray-PE and its extended form, Log-PE, enable Spikformer and QKFormer to achieve the best performance among their variants. CPG-PE is a spiking version of absolute PE designed for SNNs. Spikformer and QKFormer, when equipped with our proposed Gray-PE and Log-PE, consistently outperform all other corresponding variants, including CPG-PE and Binary-PE. Furthermore, we were surprised to find that Binary-PE fails to capture the relative positional information. This leads us to reconsider the underlying issue: if a binary encoding method cannot represent distances between certain positions, it is unsuitable as an RPE solution for SNNs. In contrast, Gray Code and Log-PE provide approximate binary encoding approaches that “partially” achieve this goal.
(4) For long input sequences, Log-PE is more effective than Gray-PE in capturing relative positional information. The input sequence length for Metr-la and Pems-bay is $12$, whereas, for Solar and Electricity, it is $168$. As an approximated approach of RPE, Gray-PE can only represent the distance between two indexes differing by $2^{n}$. This limitation means that SNNs using Gray-PE may experience a performance drop when processing longer input sequences. Furthermore, on the long-sequence datasets Solar and Electricity, SNNs equipped with Log-PE consistently outperform those with Gray-PE across nearly all prediction length settings. For instance, on the Solar dataset with a prediction length of $48$, QKFormer with Log-PE significantly outperforms QKFormer with Gray-PE, showing an average improvement of $0.038$ in R2. This result indicates that Log-PE is more effective for processing long input sequences.
5.3 Text Classification
Text classification is another representative task for which RPE plays an important role. Therefore, we conduct experiments to evaluate the performance of SNNs with RPE in text classification tasks, comparing it against other PE techniques to highlight its advantages in capturing complex linguistic structures and contextual nuances. For this part, we strictly follow the experimental setup used in the text classification experiments in Lv et al. (2024a), and the results are presented in Table 2.
Based on the results in Table 2, it is evident that our proposed Gray-PE and Log-PE significantly outperform the other spiking positional encoding methods across several key benchmarks. Both Gray-PE and Log-PE demonstrate superior accuracy on the English and Chinese datasets, with particularly notable improvements on MR, SST-2, Subj, and ChnSenti. However, the performance of RPE on the Waimai dataset is not as strong as that of CPG-PE. We attribute this to the nature of the dataset, which consists of user reviews often containing informal language, typos, or mixed expressions. This noise can hinder the model’s ability to extract meaningful patterns. These results highlight the advantages of our proposed spiking RPE techniques, especially in handling the dependencies and varying word order in text classification tasks. Unlike spiking absolute PE, i.e., CPG-PE, which struggles to adapt to the nuances of language, Gray-PE and Log-PE provide a more flexible and context-sensitive representation, improving the model’s ability to classify sentences accurately.
Table 3: Accuracy (%) on image classification benchmarks. Numbers with denote our implementation, representing results that differ from those reported in the original papers. The best and second-best results are highlighted in bold and underlined formats, respectively. “R” denotes relative PE, while “A” denotes absolute PE. The results with shading are ours. Results are averaged across $4$ random seeds.
Model PE CIFAR10 CIFAR10-DVS CIFAR100 Avg.
Spike Type Param (M) Accuracy Param (M) Accuracy Param (M) Accuracy
Vision-Transformer A $9.32$ $\bm{96.73}$ $9.36$ $\bm{81.02}$
Spikformer w/o PE $8.00$ $93.98$ $1.99$ $76.60$ $8.04$ $74.02$ $81.53$
Spikformer w/ Conv-PE (Original) A $9.32$   $\;\;94.80^{*}$ $2.57$   $\;\;78.10^{*}$ $9.36$   $\;\;77.04^{*}$ $83.31$
Spikformer w/ CPG-PE A $8.17$ $95.06$ $2.06$ $\underline{78.40}$ $8.20$ $77.82$ $83.76$
Spikformer-XNOR w/o PE $8.00$ $94.14$ $1.99$ $76.70$ $8.04$ $74.16$ $81.67$
Spikformer-XNOR w/ Conv-PE A $9.32$ $95.20$ $2.57$ $77.90$ $9.36$ $78.04$ $83.71$
Spikformer-XNOR w/ Gray-PE 1D R $8.00$ $\underline{95.46}$ $1.99$ $77.90$ $8.04$ $\underline{78.12}$ $\underline{83.83}$
Spikformer-XNOR w/ Gray-PE 2D R $8.00$ $\bf 95.66$ $1.99$ $\bf 78.70$ $8.04$ $\bf 78.45$ $\bf 84.27$
QKFormer w/o PE (Original) $6.74$ $\;\;\underline{96.32}^{*}$ $1.50$   $\;\;\bf 83.40^{*}$ $6.76$   $\;\;\bf 80.90^{*}$ $\bf 86.87$
QKFormer w/ CPG-PE A $7.01$ $96.30$ $1.58$ $82.00$ $7.04$ $80.52$ $86.27$
QKFormer-XNOR w/o PE $6.74$ $96.28$ $1.50$ $82.90$ $6.76$ $80.80$ $86.66$
QKFormer-XNOR w/ Gray-PE 1D R $6.74$ $96.22$ $1.50$ $82.20$ $6.76$ $80.48$ $86.30$
QKFormer-XNOR w/ Gray-PE 2D R $6.74$ $\bf 96.36$ $1.50$ $\underline{83.10}$ $6.76$ $\underline{80.82}$ $\underline{86.76}$
5.4 Patch-based Image Classification
We also conduct experiments on ViT-based SNNs, including Spikformer and the current SOTA model, QKFormer. As discussed in Section 4.5, we extend Gray-PE into a two-dimensional form and integrate it into the patch-based framework to enhance the spatial relationship understanding of spiking Transformers. The results of these experiments are presented in Table 3, where we draw conclusions that:
(1) Gray-PE enhances the performance of Spikformer with less parameters. Gray-PE, whether in 1D or 2D form, consistently improves Spikformer’s classification accuracy. As an approximation of RPE, Gray-PE successfully outperforms spiking absolute PE (i.e., CPG-PE), highlighting the strong dependency between patches within an image. Notably, Spikformer-XNOR with Gray-PE 2D achieves the highest performance across all three datasets, demonstrating the effectiveness of 2D form Gray-PE in capturing relative positional information.
(2) On image classification, QKFormer is insensitive to positional encoding. Surprisingly, QKFormer shows limited improvements or even performance drops when using PE techniques. Whether with CPG-PE or Gray-PE, QKFormer fails to achieve significant performance gains and, in fact, shows a decline in performance. We attribute this phenomenon to the special attention mechanism in QKFormer, known as “q-k attention”, which sums the query matrix along the channel dimension before performing a dot product with key matrix. This strategy eliminates the positional representation provided by both CPG-PE and Gray-PE, which negatively impacts performance.
(3) For Gray-PE, the 2D form performs better than the 1D form on processing image patches. Through a comparison of the performance of SNNs using Gray-PE 1D and 2D, we find that the design of the two-dimensional form of Gray Code is effective. In the case of Spikformers, Gray-PE 2D significantly outperforms Gray-PE 1D, with a $0.44\%$ accuracy improvement averagely. Even for QKFormer, where PE’s influence is limited, Gray-PE 2D still far outperforms its 1D variant. These results confirm that Gray-PE 2D is superior to Gray-PE 1D in handling positional information for image classification tasks, validating the success of our proposed Gray-PE 2D.
5.5 Positional Encoding Analysis
In this section, we will analyze the following points: (1) Impact of the number of bits on the performance of SNNs when using Gray-PE. (2) Rationale behind why we use an approximate RPE rather than a complete one in SNNs.
Figure 3: Spikformer-XNOR with Gray-PE across various bit numbers ranging from $2$ to $12$ on (a) time-series forecasting tasks and (b) text classification tasks.
For the first question, we present the results of our parameter sensitivity experiments in Figure 3. Consider that: if the number of bits used for encoding relative positions in Gray-PE is $b$, then the total number of unique encodings possible is $2^{b}$. We set the maximum sequence length is $L$, so relative distances range from $1$ to $L-1$. According to the pigeonhole principle, if $L-1>2^{b}$, there will inevitably be at least two distances that are represented identically. This issue can be mitigated by increasing $b$ to cover the range of relative distances up to $L-1$. From Figure 3 (a), we observe that for long-sequence datasets, such as Solar and Electricity (Length $=168$), the number of bits should be at least $7$ to avoid Gray-PE missing relative positional information. However, for shorter datasets like Metr-la (Length $=12$) and ChnSenti (Length $=32$), $5$ bits are sufficient.
Table 4: Time-series forecasting performance of SNNs trained with Complete RPE (C-RPE) method. We take Metr-la and Electricity as examples. denotes failure to converge.
Model (Prediction Length $=24$) Metr-la ($L=12$) Electricity ($L=168$)
R2 $\uparrow$ RSE $\downarrow$ R2 $\uparrow$ RSE $\downarrow$
Spikformer-XNOR w/o PE .$491$$753$$955$375.375.375.375
Spikformer-XNOR w/ Log-PE .$\bf.535$$\bf.719$$\bf.974$$\bf.300$
Spikformer-XNOR w/ C-RPE .$158^{*}.$$967^{*}.$$710^{*}.$ $1.03^{*}$
QKFormer-XNOR w/o PE .$501$$740$$963$352.352.352.352
QKFormer-XNOR w/ Log-PE .$\bf.541$$\bf.715$$\bf.974$$\bf.285$
QKFormer-XNOR w/ C-RPE .$430$$824$$710^{*}.$ $1.03^{*}$
Regarding the second question, we show the performance of SNNs using complete RPE (C-RPE). As discussed in Section 4.4, C-RPE is obtained by setting $R_{i,j}$ as $\frac{L-1}{|i-j|}$ and adding it to the attention map. However, the large values of this method can cause the vanilla attention map to fail in recording accurate similarity scores. As shown in Table 4, we observe that SNNs with C-RPE significantly underperform compared to other variants, with some even failing to converge, due to the detrimental impact of the large positional encoding values on the training process.
Besides, we design another extended form of Gray-PE, leveraging hierarchical relative positional representation. However, we find that when an encoding method involves carry-over, ensuring the uniqueness of the spike representation for each relative distance becomes challenging. A detailed analysis of this issue can be found in Appendix C.
6 Conclusion
In this paper, we propose a series of relative positional encoding methods for spiking Transformers. Our approach preserves the spiking nature of SNNs while effectively representing relative distances. Experimental evaluations on time series forecasting, text classification, and image classification demonstrate significant performance improvements. These results validate the effectiveness of the proposed RPE method and highlight its potential to enhance the versatility and applicability of SNNs across various domains. Future work and limitations are discussed in Appendix D.
Impact Statement
This research enhances spiking neural networks by introducing an approximation method for relative positional encoding method based on Gray Code. The proposed approach broadens the applicability of SNNs across diverse domains, enabling more robust and versatile neural architectures. Lastly, we do not think our work will negatively impact ethical aspects or future societal consequences.
Appendix A Proof of Theorem 4.1
This section provides a detailed mathematical proof of Theorem 4.1.
A.1 Statement
For any non-negative integer $n$, and for any pair of integers $a$ and $b=a+2^{n}$, the Hamming distance between their Gray Code representations $G(a)$ and $G(b)$ is:
$d_{H}(G(a),G(b))=\begin{cases}1&\text{if }n=0,\\ 2&\text{if }n\geq 1.\end{cases}$ (17)
A.2 Proof
We aim to determine the Hamming distance $d_{H}(G(a),G(a+2^{n}))$.
A.2.1 Case 1: $n=0$
When $n=0$, $b=a+1$. By the definition of the Gray Code:
$G(a)=a\oplus\left(a\gg 1\right),\quad G(a+1)=(a+1)\oplus\left((a+1)\gg 1\right).$ (18)
The Gray Code inherently ensures that consecutive integers differ by exactly one bit. Therefore:
$d_{H}(G(a),G(a+1))=1.$ (19)
This establishes the base case for $n=0$.
A.2.2 Case 2: $n\geq 1$
For $n\geq 1$, consider $b=a+2^{n}$. We analyze the XOR of their Gray Codes:
$\begin{split}&G(a)\oplus G(a+2^{n})\\ =&\left(a\oplus\left(a\gg 1\right)\right)\oplus\left((a+2^{n})\oplus\left((a+2^{n})\gg 1\right)\right).\end{split}$ (20)
Rearranging the terms:
$\begin{split}&G(a)\oplus G(a+2^{n})\\ =&\left(a\oplus(a+2^{n})\right)\oplus\left((a\gg 1)\oplus\left((a+2^{n})\gg 1\right)\right).\end{split}$ (21)
Analyzing $a\oplus(a+2^{n})$: Adding $2^{n}$ to $a$ flips the $n$-th bit of $a$. Therefore:
$a\oplus(a+2^{n})=2^{n}.$ (22)
Analyzing $(a\gg 1)\oplus((a+2^{n})\gg 1)$: Shifting both $a$ and $a+2^{n}$ to the right by one bit:
$(a+2^{n})\gg 1=\left(a\gg 1\right)+2^{n-1}.$ (23)
Thus:
$\begin{split}&(a\gg 1)\oplus\left((a+2^{n})\gg 1\right)\\ =&\left(a\gg 1\right)\oplus\left(\left(a\gg 1\right)+2^{n-1}\right)\\ =&2^{n-1}.\end{split}$ (24)
Combining the Results:
$G(a)\oplus G(a+2^{n})=2^{n}\oplus 2^{n-1}.$ (25)
In binary representation, $2^{n}$ and $2^{n-1}$ have single 1’s in distinct positions. Their XOR results in a binary number with exactly two 1’s:
$2^{n}\oplus 2^{n-1}=2^{n}+2^{n-1}.$ (26)
Therefore, the Hamming weight (number of 1’s) of $G(a)\oplus G(a+2^{n})$ is $2$.
Therefore, for $n\geq 1$, since the XOR of the Gray Codes yields a binary number with exactly two 1’s, the Hamming distance between $G(a)$ and $G(a+2^{n})$ is:
$d_{H}(G(a),G(a+2^{n}))=2.$ (27)
A.3 Final Conclusion
Combining both cases, we conclude that for any non-negative integer $n$:
$d_{H}(G(a),G(a+2^{n}))=\begin{cases}1&\text{if }n=0,\\ 2&\text{if }n\geq 1.\end{cases}$ (28)
This rigorously proves the observed property of Gray Codes concerning the Hamming distance between numbers differing by powers of two.
Appendix B Experimental Settings
B.1 Datasets
B.1.1 Time-series Forecasting
We strictly follow the dataset settings of CPG-PE (Lv et al., 2024a). The datasets we used are as follows: Metr-la (Li et al., 2017b): Average traffic speed data collected from highways in Los Angeles County. Pems-bay (Li et al., 2017b): Average traffic speed data from the Bay Area. Electricity (Lai et al., 2018): Hourly electricity consumption data in kilowatt-hours (kWh) of $321$ clients. Solar (Lai et al., 2018): Solar power production. The detailed statistical characteristics and distribution ratios for each dataset are presented below:
Table 5: The statistics of time-series datasets.
Dataset Samples Variables Observation Length Train-Valid-Test Ratio
Metr-la $34,272$ $207$ $12,(\operatorname{short-term})$ $(0.7,0.2,0.1)$
Pems-bay $52,116$ $325$ $12,(\operatorname{short-term})$ $(0.7,0.2,0.1)$
Solar-energy $52,560$ $137$ $168,(\operatorname{long-term})$ $(0.6,0.2,0.2)$
Electricity $26,304$ $321$ $168,(\operatorname{long-term})$ $(0.6,0.2,0.2)$
B.1.2 Text Classification
For text classification, we follow Lv et al. (2023b) to conduct experiments on six easy discrimination tasks, covering both English and Chinese datasets. Here are the datasets we used in text classification experiments:
The MR dataset, which stands for Movie Review, contains movie-review documents labeled based on their overall sentiment polarity (positive or negative) or subjective rating (Pang & Lee, 2005).
SST-$5$ includes $11,855$ sentences from movie reviews for sentiment classification across five categories: very negative, negative, neutral, positive, and very positive (Socher et al., 2013).
SST-$2$ is the binary version of SST-$5$, containing only two classes: positive and negative. The Subj dataset is designed to classify sentences as either subjective or objective [^1] https://www.cs.cornell.edu/people/pabo/movie-review-data/.
ChnSenti consists of approximately $7,000$ Chinese hotel reviews, each annotated with a positive or negative label [^2] https://raw.githubusercontent.com/SophonPlus/ChineseNlpCorpus/master/datasets/ChnSentiCorp_htl_all/ChnSentiCorp_htl_all.csv.
Waimai contains around $12,000$ Chinese user reviews from a food delivery platform, intended for binary sentiment classification (positive and negative) [^3] https://raw.githubusercontent.com/SophonPlus/ChineseNlpCorpus/master/datasets/waimai_10k/waimai_10k.csv.
B.1.3 Image Classification
Here are the datasets we used in image classification experiments:
The CIFAR dataset is one of the most widely used benchmarks for image classification, comprising a collection of $60,000$ color images, each with a resolution of $32\times 32$ pixels. These images are partitioned into $50,000$ training samples and $10,000$ test samples. The dataset includes 10 classes, each containing $6,000$ images, and spans a variety of object categories such as airplanes, cars, birds, and cats. The relatively low resolution of the images makes the dataset a challenging benchmark for evaluating model performance in small-scale image classification tasks.
The CIFAR10-DVS dataset represents a neuromorphic adaptation of the original CIFAR10 set, where static images have been converted into dynamic representations that simulate the recording capabilities of a Dynamic Vision Sensor (DVS) camera. Unlike traditional cameras, a DVS captures changes in the scene as individual events, rather than capturing full-frame images at fixed time intervals. This conversion results in a dataset that is more aligned with how biological vision systems process information. The CIFAR10-DVS dataset consists of $9,000$ training samples and $1,000$ test samples, with a higher resolution of $128\times 128$ pixels compared to the original CIFAR10. The event-driven nature of this dataset presents unique challenges in terms of processing and model adaptation, as it requires handling sparse, asynchronous event streams rather than dense, synchronous pixel data. This dataset is particularly valuable for testing models designed for neuromorphic systems and event-based vision tasks, offering a more realistic and biologically plausible approach to image classification.
B.2 Time-series Forecasting
Metrices
The metrics we used in time-series forecasting are the coefficient of determination (R2) and the Root Relative Squared Error (RSE).
$R^{2}$ $=\frac{1}{MCL}\sum_{m=1}^{M}\sum_{c=1}^{C}\sum_{l=1}^{L}\left[1-\frac{(Y^{m}_{c,l}-\hat{Y}^{m}_{c,l})^{2}}{(Y^{m}_{c,l}-\bar{Y}_{c,l})^{2}}\right],$ (29)
$\mathrm{RSE}$ $=\sqrt{\frac{\sum_{m=1}^{M}||\mathbf{Y}^{m}-\hat{\mathbf{Y}}^{m}||^{2}}{\sum_{m=1}^{M}||\mathbf{Y}^{m}-\bar{\mathbf{Y}}||^{2}}}.$ (30)
In these formulas, $M$ represents the size of the test set, $C$ denotes the number of channels, and $L$ signifies the length of the predictions. $\bar{\mathbf{Y}}$ is the average of $\mathbf{Y}^{m}$. The term $Y^{m}_{c,l}$ refers to the $l$-th future value of the $c$-th variable for the $m$-th sample, while $\bar{Y}{c,l}$ represents the mean of $Y^{m}{c,l}$ across all samples. The symbols $\hat{\mathbf{Y}}^{m}$ and $\hat{Y}_{c,l}^{m}$ are used to denote the predicted values. Compared to Mean Squared Error (MSE) or Mean Absolute Error (MAE), these metrics exhibit greater resilience to the absolute values of the datasets, making them especially valuable in time-series forecasting tasks.
Model Architecture
All SNNs take $4$ time steps for spiking neurons. We construct all Spikformer as $2$ blocks, setting the feature dimension as $256$, and the hidden feature dimension in FFN as $1024$. As for QKFormer, we set the block number as $4$, $2$ of which are QK blocks and the other $2$ are Spikformer blocks.
Training Hyper-parameters
we set the training batch size as $32$ and adopt Adam (Kingma & Ba, 2014) optimizer with a cosine scheduler of learning rate $1\times 10^{-4}$. An early stopping strategy with a tolerance of $30$ epochs is adopted. For other configurations, we honestly follow the SeqSNN framework  [^4] https://github.com/microsoft/SeqSNN proposed by Lv et al. (2024b). We conducted time-series forecasting experiments on 24G-V100 GPUs. On average, a single experiment takes about $1$ hour under the settings above.
B.3 Text Classification
Model Achirecture
All Spikformers are with $12$ encoder blocks and $768$ feature embedding dimension. We have substituted layer normalization of SpikeBERT (Lv et al., 2023a) with batch normalization in our directly-trained Spikformer models for text classification tasks.
Training Hyper-parameters
We directly trained Spikformers with arctangent surrogate gradients on all datasets. We use the BERT-Tokenizer in Huggingface [^5] https://huggingface.co/ to tokenize the sentences to token sequences. We pad all samples to the same sequence length of $256$. We conducted text classification experiments on $4$ RTX-3090 GPUs, and set the batch size as $32$, optimizer as AdamW (Loshchilov & Hutter, 2018) with weight decay of $5\times 10^{-3}$, and set a cosine scheduler of starting learning rate of $5\times 10^{-4}$. What’s more, in order to speed up the training stage, we adopt the automatic mixed precision training strategy. On average, a single experiment takes about $1.5$ hours under the settings above.
B.4 Image Classification
Model Architecture
For all Spikformer models, we standardized the configuration to include $4$ time steps. Specifically, for the CIFAR10 and CIFAR100 datasets, the models were uniformized with $4$ encoder blocks and a feature embedding dimension of $384$. For the CIFAR10-DVS dataset, the models were adjusted to have $2$ encoder blocks and a feature embedding dimension of $256$. For all QKFormers, we set the block number as $4$, where $2$ blocks are QK blocks and the other $2$ are Spikformer blocks.
Training Hyper-parameters
We honestly follow the experimental settings in Spikformer (Zhou et al., 2023b) and QKFormer (Zhou et al., 2024), whose source code and configuration files are available at https://github.com/ZK-Zhou/spikformer and https://github.com/zhouchenlin2096/QKFormer. As the training epochs are quite big ($300$ or $400$ epochs) in their settings, we choose to use one 80G-A100 GPU, and it takes about $3$ hours to conduct a single experiment, on average.
Appendix C Analysis on Carry-Over
We observe that when the encoding method involves carry-over, as in binary systems, it becomes difficult to ensure the uniqueness of the spike representation for each relative distance, especially with limited dimensions in the query and key vectors during self-attention. The carry-over operation introduces dependencies between bits, meaning the encoding of one bit can influence the encoding of higher-order bits. Consequently, the Hamming distance between the encodings of different relative distances is not fixed and can vary significantly due to carry-over. These dependencies also increase the likelihood of encoding collisions, where different relative distances might share identical or similar binary encodings. This can be empirically demonstrated by the poor performance of SNNs with Binary-PE in time-series forecasting, as shown in Table 1. From this perspective, Gray-PE and its extended variants can be viewed as approximate solutions, as they partially address the challenge of ensuring unique spike representations for a range of relative distances.
Appendix D Limitations and Future Work
D.1 Limitations
Despite the promising enhancements introduced by our Gray Code-based relative positional encoding method for spiking neural networks, several limitations must be acknowledged. Firstly, the current Gray Code implementation may encounter scalability issues when applied to very long input sequences or more complex network architectures. Additionally, while Gray Code effectively preserves binary spike representations, it may limit the flexibility and adaptability of the encoding scheme across diverse data modalities and task requirements. Furthermore, our evaluations have been confined to specific applications such as time series forecasting, text classification, and image patch classification, which may not fully capture the method’s performance in other domains, such as object detection (Luo et al., 2024) and real-world geometry representation (Liao et al., 2024).
D.2 Future Work
Future work should focus on optimizing the Gray Code-based RPE to enhance its scalability and efficiency, enabling its deployment in larger and more intricate SNN models. Exploring alternative encoding strategies or hybrid approaches could provide greater flexibility and improve the robustness of positional encoding across various data types and tasks. Expanding the scope of evaluation to include a wider range of applications would offer a more comprehensive understanding of the method’s effectiveness. Additionally, integrating Gray Code-based RPE with other advanced neural network components, such as attention mechanisms or neuromorphic hardware, could further elevate the performance and practical utility of SNNs. These efforts will contribute to the advancement of more versatile and powerful biologically inspired neural network architectures.

### v2

@@
 Abstract
-Spiking neural networks (SNNs) are bio-inspired networks that model how neurons in the brain communicate through discrete spikes, which have great potential in various tasks due to their energy efficiency and temporal processing capabilities.
-SNNs with self-attention mechanisms (Spiking Transformers) have recently shown great advancements in various tasks such as sequential modeling and image classifications.
-However, integrating positional information, which is essential for capturing sequential relationships in data, remains a challenge in Spiking Transformers.
-In this paper, we introduce an approximate method for relative positional encoding (RPE) in Spiking Transformers, leveraging Gray Code as the foundation for our approach.
-We provide comprehensive proof of the method’s effectiveness in partially capturing relative positional information for sequential tasks.
-Additionally, we extend our RPE approach by adapting it to a two-dimensional form suitable for image patch processing.
-We evaluate the proposed RPE methods on several tasks, including time series forecasting, text classification, and patch-based image classification.
-Our experimental results demonstrate that the incorporation of RPE significantly enhances performance by effectively capturing relative positional information.
-Spiking Neural Networks, Relative Positional Encoding
+Spiking neural networks (SNNs) are bio-inspired networks that mimic how neurons in the brain communicate through discrete spikes, which have great potential in various tasks due to their energy efficiency and temporal processing capabilities.
+SNNs with self-attention mechanisms (spiking Transformers) have recently shown great advancements in various tasks, and inspired by traditional Transformers, several studies have demonstrated that spiking absolute positional encoding can help capture sequential relationships for input data, enhancing the capabilities of spiking Transformers for tasks such as sequential modeling and image classification.
+However, how to incorporate relative positional information into SNNs remains a challenge.
+In this paper, we introduce several strategies to approximate relative positional encoding (RPE) in spiking Transformers while preserving the binary nature of spikes.
+Firstly, we formally prove that encoding relative distances with Gray Code ensures that the binary representations of positional indices maintain a constant Hamming distance whenever their decimal values differ by a power of two, and we propose Gray-PE based on this property.
+In addition, we propose another RPE method called Log-PE, which combines the logarithmic form of the relative distance matrix directly into the spiking attention map.
+Furthermore, we extend our RPE methods to a two-dimensional form, making them suitable for processing image patches.
+We evaluate our RPE methods on various tasks, including time series forecasting, text classification, and patch-based image classification, and the experimental results demonstrate a satisfying performance gain by incorporating our RPE methods across many architectures.
+Our results provide fresh perspectives on designing spiking Transformers to advance their sequential modeling capability, thereby expanding their applicability across various domains.
 1 Introduction
-Spiking Neural Networks (SNNs) (Maass, 1997) are a class of bio-inspired models designed to emulate the communication process of biological neurons, which transmit information through discrete spikes.
+Spiking Neural Networks (SNNs) [1] are a class of bio-inspired models designed to emulate the communication process of biological neurons, which transmit information through discrete spikes.
 In contrast to artificial neural networks (ANNs) that operate on continuous values, SNNs process information in the form of spikes occurring at precise moments in time.
-This discrete event-driven processing enables SNNs to be highly energy-efficient.
-The temporal characteristics of spikes make SNNs particularly well-suited for tasks involving sequential data or dynamic environments, such as sensory processing (Li et al., 2017a; Fang et al., 2021), patch-based image classification (Zhou et al., 2023b; Yao et al., 2023), time-series forecasting (Lv et al., 2024b; Park et al., 2024), and natural language processing (Zhu et al., 2023; Lv et al., 2023b; Xing et al., 2024).
-In ANNs, relative positional encoding (RPE) is often used to encode the positions of input elements (Su et al., 2024), enabling the model to capture the relative relationships between positions within a sequence.
-While RPE has been shown to enhance performance in language modeling (Su et al., 2024) and image processing (Liu et al., 2021), its direct application to SNNs remains limited due to the unique characteristics of spike-based computation.
-Current approaches to realize positional encoding (PE) in SNNs either fail to maintain the uniqueness of spike representations for each position (Zhou et al., 2023b; Yao et al., 2024a), or ignore the incorporation of relative positional information (Lv et al., 2024a).
-In this paper, we first identify a limitation in the current spiking self-attention (SSA) mechanism (Zhou et al., 2023b), which is suboptimal for measuring the distance (or similarity) between tokens in the query and key matrices.
-Specifically, the dot product operation fails to effectively differentiate distances when both the query and key are spike matrices.
-To address this limitation, we propose a new spiking self-attention mechanism based on the Exclusive-NOR (XNOR) operation, which enhances the measurement of query-key similarity by leveraging the Hamming distance (see Section 3.4).
-Secondly, we introduce a novel approximate relative positional encoding method for SNNs, based on Gray Code (Gray, 1953) as the foundational encoding scheme.
-The consistent Hamming distance between Gray Code representations of two position indices differing by $2^{n}$ allows for serving as a spiking RPE method and the integration with the XNOR-based SSA mechanism.
-The binary encoding nature of Gray Code is well-suited to represent relative positions while preserving the spiking dynamics of SNNs, which is crucial for effectively processing sequential data.
-Furthermore, we provide a theoretical proof in Appendix A demonstrating the efficacy of Gray Code in partially capturing relative positional information for sequential tasks.
-Moreover, we extend our RPE method to the logarithmic form that can be directly integrated into the attention map of XNOR-based spiking self-attention, offering greater efficiency compared to the previous form while maintaining comparable performance.
-For image patches, we also design a two-dimensional version of our method to capture the relative positional information of both height and width.
-To evaluate the effectiveness of our RPE approaches, we apply them to three diverse tasks: time series forecasting, text classification, and patch-based image classification.
-Our experiments demonstrate that the proposed RPE method enhances the performance of SNNs across all three tasks, allowing the original SNNs to capture and leverage relative relationships between elements in the input sequential data.
-Moreover, we further conduct an ablation study and positional encoding analysis to validate the inner properties of our proposed RPE method.
-Based on the analysis, we offer several valuable insights into designing effective positional encoding schemes for SNNs.
-We believe that this work contributes a new method for integrating relative positional encoding into SNNs, expanding their potential for a wide range of applications in machine learning and computational neuroscience.
-To summarize, our contributions are as follows:
-  • A Novel RPE Method for SNNs.
-Our research represents one of the pioneering efforts to propose a solution to relative positional encoding for spiking neural networks, albeit as an approximation method.
-  • Consistent Performance Improvement.
-Our proposed method and its extended variants significantly and consistently improve the performance of SNNs across various sequential tasks, including time-series forecasting and text classification.
-  • Theoretical Validation and Insightful Analysis.
-In addition to experimental evaluations, we provide thorough theoretical proof to validate the effectiveness of our RPE method.
-Furthermore, we offer an analysis of the underlying properties of RPE, providing valuable insights for future designs of RPE in SNNs.
+The temporal characteristics of spikes make SNNs particularly well-suited for tasks involving sequential data or dynamic environments, such as sensory processing [2, 3], patch-based image classification [4, 5], time-series forecasting [6, 7, 8], and natural language processing [9, 10, 11].
+In the vanilla Transformer architecture [12], positional encoding serves as a critical mechanism for modeling sequential dependencies in input data.
+Beyond absolute positional encoding, relative positional encoding (RPE) [13, 14] has emerged as an effective approach to represent inter-element distances, enabling models to capture relational patterns within sequences dynamically.
+Although RPE has demonstrated effectiveness in improving language modeling [13] and visual recognition tasks [15], its integration into SNNs remains underexplored.
+Existing methodologies for implementing positional encoding in spiking Transformers either suffer from ambiguous spike representations across positions [4, 16], or neglect to integrate relative positional relationships entirely [7].
+Directly adapting current RPE techniques, such as Attention with Linear Biases (ALiBi) [13] and Rotary Position Embedding (RoPE) [14], to spiking Transformers encounters significant challenges.
+Specifically, spiking neural architectures exhibit intrinsic difficulty in decoupling relative positional information from their sparse, event-driven representations.
+This limitation, empirically demonstrated in Section 5.2, underscores the necessity for rethinking RPE integration to align with neuromorphic computing principles, such as temporal sparsity and spike-based communication.
+In this paper, we first propose that the Hamming distance [17], which quantifies the number of ones resulting from the XOR operation between two binary strings, serves as an appropriate metric for measuring relative distances when both the query and key matrices are binary.
+Consequently, we refine the spiking self-attention mechanism [4] by replacing dot-product operations with exclusive-NOR (XNOR) logic operations.
+Then we present two novel approximation strategies for integrating RPE into spiking Transformers, while strictly preserving the binary activation dynamics inherent to spiking neurons.
+First, we propose Gray-PE, a method exploiting the properties of Gray Code [18] to binarize relative positional distances.
+We theoretically prove that encoding relative distances via Gray Code ensures a constant Hamming distance between the binary representations of positional indices whose decimal differences equal $2^{n}$, where $n\geq 0$ (See Theorem 1).
+This property guarantees that any pair of positions separated by a relative distance of $2^{n}$ in decimal space exhibits invariant Hamming distances in their Gray Code-encoded representations.
+Such invariance stabilizes positional relationship modeling for power-of-two intervals, addressing a critical limitation in existing spiking neural architectures.
+Second, we propose Log-PE, a method adapting insights from ALiBi [13] and Rectified RoPE [19].
+Log-PE integrates a non-negative logarithmic transformation of the relative distance map directly into the spiking attention map, inducing a decaying sensitivity to positional relationships akin to windowed attention mechanisms.
+Moreover, we extend the proposed RPE methods to their two-dimensional form, making them suitable for processing image patches.
+To systematically evaluate the efficacy of our proposed RPE methods, we benchmark them across three cross-domain tasks: time series forecasting, text classification, and patch-based image classification.
+We employ three representative spiking Transformer architectures as backbones: Spikformer [4], the Spike-driven Transformer [5], and QKFormer [20].
+Experimental results demonstrate consistent performance gains across all tasks when integrating our RPE approaches, affirming that explicit modeling of relative positional relationships addresses a critical limitation in existing spiking Transformer designs.
+Furthermore, we conduct experiments on ablation study, long sequence modeling, and sensitivity analysis to validate the inner properties of our proposed RPE method.
+This work establishes a framework for integrating relative positional encoding (RPE) into spiking Transformers, advancing their applicability in neuro-inspired machine learning paradigms.
+Our primary contributions are summarized as follows:
+  • Two RPE Methods for Spiking Transformers.
+To our knowledge, this study is among the first to explore RPE adaptations for spiking architectures systematically.
+While Gray-PE and Log-PE operate as principled approximations constrained by binary spike dynamics, they address a critical gap in positional modeling for neuromorphic computation.
+  • Theoretical Foundations and Empirical Analysis.
+In addition to empirical validation, we provide theoretical guarantees demonstrating that our methods can partially encode relative positional information.
+Furthermore, we offer necessary analysis on the internal properties of RPE and their robustness facing long sequences.
+  • Consistent Performance Gains Across Architectures and Tasks.
+Our proposed RPE methods consistently improve the performance of spiking Transformers across various sequential tasks, including time-series forecasting and text classification.
 2 Related Work
-In sequential tasks, positional encoding (PE) plays a crucial role in capturing the order of tokens.
-Traditional absolute positional encoding assigns a fixed position to each token, whereas relative positional encoding (RPE) emphasizes the distance between tokens.
+Positional encoding serves as an indispensable mechanism for preserving the order of input elements in sequential modeling tasks.
+Traditional absolute positional encoding assigns static, predefined embeddings to individual tokens based on their sequential indices.
+In contrast, relative positional encoding (RPE) dynamically models the pairwise distances between tokens, enabling the self-attention mechanism to prioritize interactions based on their relative proximity.
 RPE allows the model to generalize across different sequence lengths and better capture relationships between tokens.
-Despite its significance, there has been limited research focusing on the role of positional encoding approaches in spiking neural networks.
-Notably, both Spikformer (Zhou et al., 2023b) and Spike-driven Transformer (Yao et al., 2023, 2024a) incorporate a combination of convolutional layers, batch normalization, and spiking neuron layers to generate “learnable relative positional encodings”.
-However, we argue that this approach functions more similarly to a spike-element-wise residual connection (Fang et al., 2021) than to a conventional positional encoding module.
-A robust PE module should offer unique representations for each position, but the spike-position matrices generated by these methods may lead to identical spike representations for different positions.
-Lv et al.
-(2024a) introduced a spiking positional encoding method for SNNs, inspired by central pattern generators (Marder & Bucher, 2001).
-This method generates periodic positional encoding, ensuring a unique binary spike matrix for each position.
+Despite the importance of PE in sequence-aware architectures, its application to SNNs is limited.
+Existing implementations, such as Spikformer [4] and Spike-driven Transformer [5, 16, 21], incorporate a combination of convolutional layers, batch normalization, and spiking neuron layers to derive learnable positional encodings.
+However, we argue that this approach functions more similarly to a spike-element-wise residual connection [3] than to a conventional positional encoding module.
+A principled PE module should offer unique representations for positions, but the spike-position matrices generated by these methods may lead to identical spike representations for different positions.
+CPG-PE, proposed by [7], introduces a spiking absolute positional encoding inspired by central pattern generators [22], generating unique periodic binary spike patterns for each position.
 However, their approach is based on absolute positional encoding and, thus, does not capture the time-translational invariance property in many sequential modeling problems, which, however, is an important advantage of relative positional encoding methods.
 3 Preliminary
-Figure 1: Preliminary Figures.
-(a) Spike dynamics of spiking (LIF) neurons.
-(b) Illustration of spiking self-attention in Spikformer (Zhou et al., 2023b) and QKFormer (Zhou et al., 2024).
-(c) An example of Hamming Distance.
-(d) Calculation process of Gray Code.
-3.1 Spiking Neural Networks
-The fundamental building block of SNNs is the spiking neuron, such as the leaky integrate-and-fire (LIF) model (Maass, 1997), which is governed by the input current $I(t)$, influencing the membrane potential $U(t)$ and the spike output $S(t)$ at each time step $t$.
-The dynamics of the LIF neuron are captured by the following system of equations:
-$U(t)=H(t-\Delta t)+I(t),\quad I(t)=f(\mathbf{x};\mathbf{\theta}),$ (1)
-$H(t)=V_{\rm reset}S(t)+(1-S(t))\beta U(t),$ (2)
-$S(t)=\begin{cases}1,&\text{if }U(t)\geq U_{\rm thr},\\ 0,&\text{if }U(t)<U_{\rm thr}.\end{cases}$ (3)
-Here, $I(t)$ represents the spatial input to the LIF neuron at time $t$, which is computed by the function $f$ using the input vector $\mathbf{x}$ and the learnable parameters $\mathbf{\theta}$.
-The discretization constant $\Delta t$ controls the time granularity in the LIF model, and $H(t)$ is the temporal output at time step $t$.
-The spike $S(t)$ is determined by a Heaviside step function that is activated when the membrane potential $U(t)$ exceeds a threshold $U_{\rm thr}$.
-Upon reaching the threshold, the neuron generates a spike, and $H(t)$ resets to $V_{\rm reset}$.
-If $U(t)$ does not exceed the threshold, the membrane potential decays to $H(t)$ at a rate determined by $\beta$.
+Figure 1: Illustration of preliminary knowledge.
+(a) Spike dynamics of LIF neurons.
+(b) Illustration of vanilla spiking self-attention in Spikformer [4].
+(c) An example of Hamming Distance between two spike trains.
+(d) The calculation process of the classic Reflected Gray Code.
+3.1 Spiking Neurons
+We take the leaky integrate-and-fire (LIF) neuron [1] as our building brick of SNNs, which is governed by the input current $I[t]$, influencing the membrane potential $U[t]$ and the spike output $S[t]$ at each time step $t$.
+The dynamic of the LIF neuron is captured by the following system of equations:
+$U[t]$ $=H[t](1-S[t])+U_{\text{reset}}S[t],\quad S[t]=\Theta(H[t]-U_{\text{thr}}),$ (1)
+$H[t]$ $=U[t-1]+\frac{1}{\tau}(I[t]-(U[t-1]-U_{\text{reset}})),$ (2)
+where $\tau$ is the membrane time constant.
+The spike $S(t)$ will be triggered when the membrane potential $H(t)$ exceeds a threshold $U_{\text{thr}}$, right after which $U[t]$ will be reset to $U_{\text{reset}}$.
 3.2 Spiking Self-Attention
-Spiking self-attention (SSA) is a spiking version of self-attention (Vaswani, 2017), which was proposed in Spikformer (Zhou et al., 2023b) and QKFormer (Zhou et al., 2024).
+Spiking self-attention (SSA) is a spiking version of self-attention [12], which was proposed in Spikformer [4].
 The vital design is to utilize discrete spikes to approximate the vanilla self-attention mechanism.
 It can be written as:
-$\small\begin{split}\mathbf{Q}&=\mathcal{SN}\left(\operatorname{BN}\left(\mathbf{X}\cdot\mathbf{W_{Q}}\right)\right),\\ \mathbf{K}&=\mathcal{SN}\left(\operatorname{BN}\left(\mathbf{X}\cdot\mathbf{W_{K}}\right)\right),\\ \mathbf{V}&=\mathcal{SN}\left(\operatorname{BN}\left(\mathbf{X}\cdot\mathbf{W_{V}}\right)\right),\end{split}$ (4)
+$\mathbf{Q},\mathbf{K},\mathbf{V}$ $=\mathcal{SN}\left(\operatorname{BN}\left(\mathbf{X}\cdot\mathbf{W}_{{Q,K,V}}\right)\right)\in\{0,1\}^{T\times L\times D}$ (3)
 where $\mathcal{SN}$ is a spike neuron layer described in Equation 1.
 The input is denoted as $\mathbf{X}\in\{0,1\}^{T\times L\times D}$, where $T$ is the number of time steps.
-$\operatorname{BN}$ represents batch normalization, and $\tau$ is a scaling factor.
+$\operatorname{BN}$ represents batch normalization, and $\sigma$ is a scaling factor.
 The attention map $\mathbf{AttnMap}$ is then computed as the dot product between $\mathbf{Q}$ and $\mathbf{K}^{\operatorname{T}}$:
-$\small\mathbf{AttnMap}=\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}},$ (5)
-$\small\operatorname{SSA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)=\mathcal{SN}(\operatorname{BN}((\mathbf{Attn}\cdot\mathbf{V}*\tau)\cdot\mathbf{W})).$ (6)
+$\operatorname{SSA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)=\mathcal{SN}(\operatorname{BN}((\underbrace{\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}}}_{\mathbf{AttnMap}}\cdot\mathbf{V}*\sigma)\cdot\mathbf{W})).$ (4)
 As a result, the attention map $\mathbf{AttnMap}\in\mathbb{N}_{0}^{T\times L\times L}$, where $\mathbb{N}_{0}$ denotes the set of non-negative integers.
@@
 The parameters $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, $\mathbf{W}_{V}$, and $\mathbf{W}$ are all learnable parameters.
+Recent studies, including Spike-Driven Transformer (SDT) [5, 16, 21], SpikingResFormer [23], and QKFormer [20], have proposed various modifications to the standard SSA mechanism.
+For our empirical evaluation, we selectively employ architectures demonstrating compatibility with our proposed relative position encoding methods.
 3.3 Relative Positional Encoding
-In self-attention mechanisms, the attention weights are modified using a relative positional encoding (RPE) $\mathbf{R}_{i,j}$, which depends on the distance between the $i$-th and $j$-th tokens.
-A typical realization of RPE (Liu et al., 2021; Hao et al., 2024) is formulated as:
-$\small\begin{split}\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})&=\mathbf{AttnMap\cdot V}\\ &=\text{Softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}}}{\sqrt{d_{k}}}+\mathbf{R}_{i,j}\right)\mathbf{V}.\end{split}$ (7)
-Here, $\mathbf{R}_{i,j}$ represents the relative distance between tokens $i$ and $j$.
-Specifically, the relative positional encoding for the $i$-th position in $\mathbf{Q}$ and the $j$-th position in $\mathbf{K}$ should satisfy distance consistency, i.e., the magnitude of $\mathbf{R}_{i,j}$ should depend solely on the relative distance $|i-j|$, enabling the model to distinguish between nearby and distant tokens.
-This approach improves the model’s ability to handle long-range dependencies and sequence variations.
-Figure 2: Overview of Our Method.
-(a) XNOR-based Spiking Self-Attention (SSA).
-This panel illustrates the computation flow for the query ($\mathbf{Q}$) and key ($\mathbf{K}$) matrices in a PyTorch-style notation.
-(b) Gray-PE.
-Gray code representations exhibit a consistent Hamming distance for two integers differing by $2^{n}$.
-Gray-PE is implemented by concatenating $G(\bm{l})$ along the $D$ dimension of both $\mathbf{Q}$ and $\mathbf{K}$.
-(c) Log-PE.
-A pre-assigned relative distance encoding map $\mathbf{R}_{i,j}\in\mathbb{N}_{0}$ is added to the standard attention map $\mathbf{AttnMap}$.
-The shading indicates the magnitude of the values, with deeper blue corresponding to larger values.
-(d) 2D Form of Gray-PE.
-For image patches, a 2D relative positional encoding is more suitable than the 1D version, as it better captures the spatial relationships.
+Relative positional encoding (RPE) in Transformers primarily introduces bias terms into the self-attention mechanism that dynamically encode pairwise token distances.
+A common implementation of RPE, as demonstrated in prior work [15, 24], is formalized as follows:
+$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\underbrace{\text{Softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}}}{\sqrt{d_{k}}}+\mathbf{R}_{i,j}\right)}_{\mathbf{AttnMap}}\cdot\mathbf{V}.$ (5)
+Here, $\mathbf{R}_{i,j}$ represents the relative positional bias between the $i$-th query and the $j$-th key positions.
+A critical aspect of RPE is its adherence to distance consistency: the magnitude of $\mathbf{R}_{i,j}$ is determined exclusively by the relative positional offset $|i-j|$, ensuring that the model systematically differentiates between proximal and distant tokens.
+This property enhances the model’s capacity to capture long-range dependencies and generalize across variations in sequence length and structure.
 3.4 Hamming Distance
-The Hamming distance between two binary strings of equal length is the number of bit positions at which the corresponding bits are different.
+The Hamming distance [17] between two binary strings of equal length is the number of bit positions at which the corresponding bits are different.
 Formally, for two binary strings $A$ and $B$ of length $m$,
-$\small d_{H}(A,B)=\sum_{i=1}^{m}\delta(A_{i},B_{i}),$ (8)
-where
-$\small\delta(A_{i},B_{i})=\begin{cases}1&\text{if }A_{i}\neq B_{i},\\ 0&\text{otherwise}.\end{cases}$ (9)
+$\small d_{H}(A,B)=\sum_{i=1}^{m}\delta(A_{i},B_{i}),\quad\text{where}\quad\delta(A_{i},B_{i})=\begin{cases}1&\text{if }A_{i}\neq B_{i},\\ 0&\text{otherwise}.\end{cases}$ (6)
+Hamming distance is suitable for measuring the relative distances when $\mathbf{Q}$ and $\mathbf{K}$ are spike matrices.
+3.5 Gray Code
+Gray Codes [18], also known as reflected binary codes, are binary numbering systems where adjacent values differ by precisely one bit.
+For a non-negative integer $x$, the standard binary reflected Gray Code $G(x)$ is defined by the following bitwise operation:
+$\small G(x)=x\oplus\left(x\gg 1\right),$ (7)
+where $\oplus$ denotes the bitwise XOR operation, and $\gg$ denotes the arithmetic right shift.
 Since the preliminary knowledge involved is extensive and loosely connected, we have provided Figure 1 to help readers visually grasp the key concepts of each section.
 4 Method
-4.1 Gray Code Definition
-Gray Codes (Gray, 1953), also known as reflected binary codes, are binary sequences where consecutive numbers differ by exactly one bit.
-Gray Code $G(x)$ for a non-negative integer $x$ can be generated using the formula: The standard binary reflected Gray Code $G(x)$ for a non-negative integer $x$ is defined as:
-$\small G(x)=x\oplus\left(x\gg 1\right),$ (10)
-where $\oplus$ denotes the bitwise XOR operation, and $\gg$ denotes the arithmetic right shift (with zero padding).
+Figure 2: Overview of Our Method.
+(a) XNOR-based spiking self-attention.
+We illustrate the computation flow for $\mathbf{Q}$ and $\mathbf{K}$ in a PyTorch-style notation.
+(b) Gray-PE.
+Position indices differing by $2^{n}$ exhibit a consistent Hamming distance on their Gray code representations.
+Gray-PE is implemented by concatenating $G(\bm{l})$ along the $D$ dimension on both $\mathbf{Q}$ and $\mathbf{K}$.
+(c) Log-PE.
+A pre-assigned relative distance encoding map $\mathbf{R}_{i,j}\in\mathbb{N}_{0}$ is added to the original attention map $\mathbf{AttnMap}$.
+(d) 2D Form of Gray-PE.
+A 2D RPE is more suitable than the 1D version for image patches, as it captures the spatial relationships more effectively.
+4.1 Design Principles
+Relative position encoding (RPE) aims to encode the relative distances between positional indices within a sequence.
+In many spiking Transformers, such as Spikformer, Spike-Driven Transformer, and QKFormer, both the $\mathbf{Q}$ and $\mathbf{K}$ matrices are binary.
+Consequently, their relative distances can be computed using the Hamming distance, which corresponds to the number of ones resulting from the XOR operation between $\mathbf{Q}$ and $\mathbf{K}$.
+To better align with this Hamming distance-based similarity measure, we replace the traditional dot-product spiking self-attention (SSA) mechanism with an XNOR-based SSA.
+Inspired by RPE strategies in Transformers, we propose two approaches for incorporating relative distance information into spiking attention mechanisms: (1) Gray-PE: Gray-Code-based positional encoding concatenated to $\mathbf{Q}$ and $\mathbf{K}$, and (2) Log-PE: logarithmic positional encoding applied directly to the attention map.
 4.2 XNOR-Based Spiking Self-Attention
-In the standard self-attention mechanism of Transformer (Vaswani, 2017), the attention map is computed as the dot product between the query matrix $\mathbf{Q}$ and key matrix $\mathbf{K}$, i.e., $\mathbf{AttnMap}=\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}}$.
-The primary concept behind this approach is to use the dot product to measure the similarity between each token in $\mathbf{Q}$ and $\mathbf{K}$, which works effectively when both $\mathbf{Q}$ and $\mathbf{K}$ are continuous-valued matrices.
-However, when $\mathbf{Q}$ and $\mathbf{K}$ are spike matrices, the dot product spike matrices, the dot product becomes inadequate for capturing the temporal or event-based interactions intrinsic to spike-based representations.
-As a result, the dot product operation in spiking self-attention, as outlined in Section 3.2, is not well-suited for spiking Transformers.
-To address this limitation, several researchers (Yao et al., 2023; Zhou et al., 2023a; Yao et al., 2024b) have proposed alternative attention mechanisms specifically designed for SNNs.
-In this paper, we propose an XNOR-based spiking self-attention mechanism, which leverages the Hamming distance (See Section 3.4) to measure the similarity between each token in $\mathbf{Q}$ and $\mathbf{K}$.
-This approach enables the integration of the Gray-code-based relative positional encoding method into the attention mechanism, allowing for more effective capture of positional information in spike-based representations.
+In the original Transformer [12], the attention map is computed via the dot product between the query and key matrices, $\mathbf{AttnMap}=\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}}$, which effectively captures similarity of $\mathbf{Q}$ and $\mathbf{K}$.
+As mentioned above, in order to capture the relative distances of spiking matrices while effectively measuring the similarity, we design the XNOR-based SSA.
+Unlike the dot-product operation, XNOR accounts for both spiking state ($1$) and the resting state (00).
 Formally, we modify Equation 4 as follows:
-$\small\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot(\mathbf{Q}\oplus\mathbf{K}),$ (11)
+$\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot(\mathbf{Q}\oplus\mathbf{K}),$ (8)
 where $\lnot$ denotes the Not operation, $\oplus$ denotes the XOR operation, and $D$ represents the channel dimension.
-It is important to note that every token in $\mathbf{Q}$ will perform XOR $\oplus$ with every token in $\mathbf{K}$, so we sum over the channel dimension $D$ to get $\mathbf{AttnMap}\in\mathbb{N}_{0}^{T\times L\times L}$, shown in Figure 2 (a).
-Compared to the dot product, the XNOR operation significantly increases the values in $\mathbf{AttnMap}\in\mathbb{N}_{0}$.
-Therefore, the scale factor $\tau$ in Equation 6 should be set to a smaller value or treated as a learnable parameter, ensuring that the firing rate of $\mathcal{SN}$ does not become excessively large.
-We will empirically demonstrate in the Experiments Section that this XNOR modification does not negatively impact the performance of the vanilla spiking self-attention.
-4.3 Approximate RPE Method Based on Gray Code
-We modify the SSA module by replacing the dot product operation with the XNOR operation, enabling the use of Hamming distance to measure the similarity between tokens.
-In this context, we propose that Gray Code can serve as an approximate approach to relative positional encoding for SNNs.
-This is supported by Theorem 4.1, which demonstrates the effectiveness of Gray Code in partially capturing positional relationships within the framework of XNOR-based spiking self-attention.
-Theorem 4.1.
-For two integers differing by $2^{n}$ in decimal, their Gray Code representations have a consistent Hamming distance of $1$ for $n=0$, and $2$ for $n\geq 1$, i.e., for any non-negative integer $a$ and $n$:
-$d_{H}(G(a),G(a+2^{n}))=\begin{cases}1&\text{if }n=0,\\ 2&\text{if }n\geq 1.\end{cases}$ (12)
-Proof.
-Please refer to Appendix A for detailed proof. ∎
-From Theorem 4.1, we can infer that Gray Code ensures the consistency of distance representations for every $2^{n}$ ($n\geq 0$) relative distance.
-This property makes Gray Code an ideal approximate RPE method for SNNs, as it not only partially captures relative positions but also preserves the binary nature (00 and $1$) nature of SNNs.
-In our implementation, we concatenate the Gray Code representations of each position index to both the query matrix $\mathbf{Q}\in\mathbb{N}_{0}^{T\times L\times D}$ and key matrix $\mathbf{K}\in\mathbb{N}_{0}^{T\times L\times D}$, leaving the remaining operations unchanged.
-Formally, the attention map $\mathbf{AttnMap}$ is computed as follows:
-$\small\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot([\mathbf{Q}\parallel G(\bm{l})]\oplus[\mathbf{K}\parallel G(\bm{l})]),$ (13)
-where $G(\cdot)$ represents the function that converts non-negative integers into their corresponding Gray Code representations.
-The vector $\bm{l}$ denotes an array of position indexes, specifically $[0,1,2,\dots,L-1]$, where $L$ is the sequence length of both $\mathbf{Q}$ and $\mathbf{K}$.
+Note that every token in $\mathbf{Q}$ will perform XOR with every token in $\mathbf{K}$, so we sum over the channel dimension $D$ to get $\mathbf{AttnMap}\in\mathbb{N}_{0}^{T\times L\times L}$, shown in Figure 2 (a).
+The scale factor $\sigma$ in Equation 4 should be set to a smaller value or treated as a learnable parameter, ensuring that the firing rate of $\mathcal{SN}$ does not become excessively large.
+We will empirically demonstrate that this XNOR modification does not negatively impact the performance of the vanilla spiking self-attention.
+4.3 Gray-PE
+We propose that the Gray Code can serve as an approximate approach to relative positional encoding for spiking Transformers. This is supported by the following Theorem 1:
+Theorem 1.
+(Proof in Appendix A) For two position indices differing by $2^{n}(n\geq 0)$, their Gray Code representations have a consistent Hamming distance.
+Specifically, $\forall$ position $i$, we have: $d_{H}(G(i),G(i+2^{n}))=\begin{cases}1&\text{if }n=0,\\ 2&\text{if }n\geq 1.\end{cases}$ (9)
+As illustrated in Figure 2 (b), the Hamming distance $d_{H}(0,1)$ and $d_{H}(1,2)$ both equal $1$ because their relative distance is $1$, i.e., $02^{n},n=$ 0.
+Similarly, $d_{H}(0,2)=d_{H}(1,3)$, and $d_{H}(0,4)=d_{H}(1,5)$, as their relative distances are the power of $2$.
+That said, Gray Code ensures the consistency of relative distance representations for every $2^{n}$ ($n\geq 0$) relative distance.
+For implementation, we concatenate the Gray Code representations of each position index to both the query matrix $\mathbf{Q}\in\mathbb{N}_{0}^{T\times L\times D}$ and key matrix $\mathbf{K}\in\mathbb{N}_{0}^{T\times L\times D}$, leaving the remaining operations unchanged.
+We use concatenation instead of addition because $\mathbf{Q}$ and $\mathbf{K}$ are spike matrices, and addition would compromise their binary nature.
+Formally, the attention map $\mathbf{AttnMap}$ will be:
+$\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot([\mathbf{Q}\parallel G(\bm{l})]\oplus[\mathbf{K}\parallel G(\bm{l})]),$ (10)
+where $G(\cdot)$ represents the function that converts integers into their binary Gray Code representations.
+The vector $\bm{l}$ denotes an array of position indexes, specifically $[0,1,2,\dots,L-1]$, where $L$ is the sequence length of $\mathbf{Q}$ and $\mathbf{K}$.
 $\parallel$ denotes concatenation on the channel dimension $D$.
-We name this method as Gray-PE.
-4.4 Extended Variant on Attention Maps
-Although Gray-PE can partially capture relative distances, it faces significant challenges when the input sequence length is large or when the downstream task is highly sensitive to long-range dependencies.
-For instance, when the input sequence length reaches $10^{2}$, Gray-PE can only distinguish a limited range of relative distances.
-To address this issue, we propose an extension of Gray-PE, referred to as Log-PE.
-Specifically, we simulate Equation 7 by directly adding a pre-assigned relative position map, denoted as $\mathbf{R}_{ij}$, to the attention map produced by the XNOR-based SSA operation.
-To be formal:
-$\small\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot(\mathbf{Q}\oplus\mathbf{K})+\mathbf{R}_{i,j}$ (14)
-where
-$\mathbf{R}_{i,j}=\begin{bmatrix}R_{i,j}\end{bmatrix}=\begin{bmatrix}\lceil\log_{2}(\frac{L-1}{|i-j|})\rceil\end{bmatrix}.$ (15)
-Here, $\lceil$.$\rceil$ .
-⌉ denotes the round-up function, $L$ is the sequence length, and $i,j$ Figure 2 (c) shows an illustration of Log-PE, indicating that the farther relative distance, the smaller $R_{i,j}$.
-Since the original $\mathbf{AttnMap}\in\mathbb{N}_{0}$ is a matrix composed of non-negative integers, our primary goal is to ensure accurate relative distance consistency while preserving the effectiveness of spiking self-attention.
-If we set the $R_{i,j}$ as $\frac{L-1}{|i-j|}$, we could obtain a complete relative positional encoding for the XNOR-based SSA.
-However, we choose not to pursue this solution because, for long sequence lengths $L$, the large values of $\frac{L-1}{|i-j|}$ would severely distort the original representations in the spiking self-attention map.
-Therefore, using the logarithmic form $R_{i,j}$ represents a trade-off that balances the values of the attention map and relative positional encoding.
+Notably, the binary nature of Gray Code (comprising only 00 and $1$) aligns intrinsically with the spike-based computation paradigm, avoiding the need for floating-point operations that impose significant implementation overhead on neuromorphic hardware.
+4.4 Log-PE
+Although Gray-PE can partially capture relative distances, it faces significant challenges when the input sequence is long or when the downstream task is highly sensitive to long-range dependencies.
+For instance, when $L\geq 10^{2}$, the distinguishable range of relative distances under Gray-PE becomes constrained by its power-of-two quantization mechanism.
+To mitigate this, we propose Log-PE that integrates logarithmic positional bias into spiking-based self-attention.
+Specifically, we simulate Equation 5 and follow ALiBi [13] to directly add a pre-assigned relative position map, denoted as $\mathbf{R}_{ij}$, to the attention map produced by SSA:
+$\mathbf{AttnMap}=\left(\sum_{i=0}^{D}\lnot(\mathbf{Q}\oplus\mathbf{K})\right)+\mathbf{R}_{i,j},\;\text{where}\;\;\mathbf{R}_{i,j}=\begin{bmatrix}R_{i,j}\end{bmatrix}=\begin{bmatrix}\lceil\log_{2}(\frac{L-1}{|i-j|+1})\rceil\end{bmatrix}.$ (11)
+Here, $\lceil$.$\rceil$ . ⌉ denotes the round-up function, $L$ is the sequence length, and $i,j$ is position indices.
+Figure 2 (c) shows an illustration of Log-PE.
+Since the original $\mathbf{AttnMap}$ is a matrix composed of non-negative integers, we aim to ensure accurate relative distance consistency while preserving the effectiveness of spiking self-attention.
+Theoretically, if we set the $R_{i,j}$ as $\frac{L-1}{|i-j|+1}$, we could obtain a complete RPE for the spiking Transformers.
+However, we choose not to pursue this solution, because for long sequence lengths $L$, the large values of $\frac{L-1}{|i-j|+1}$ would catastrophically overshadow the original spiking attention activations (See Appendix B).
+Therefore, using the logarithmic form $R_{i,j}$ represents a compromise that balances the values between the spiking attention map and complete-RPE, while partially capturing relative position information.
 4.5 Two-Dimensional Form for Image Patches
-CNN-based SNN models, such as Spiking VGG (Sengupta et al., 2019) and SEW-ResNet (Fang et al., 2021), do not incorporate the concept of “positional encoding” in their spike representations.
-In contrast, Vision Transformer (Dosovitskiy et al., 2021) reformulated traditional image classification into a patch-based approach, dividing images into smaller patches.
+CNN-based SNN models, such as Spiking VGG [25] and SEW-ResNet [3], do not incorporate the concept of “positional encoding” in their spike representations.
+Vision Transformer [26] reformulated traditional image classification into a patch-based approach, dividing images into smaller patches.
 Unlike 1D positional encoding, which only considers the linear sequence of patches, 2D RPE accounts for both the horizontal and vertical positions of the patches in the image grid.
 This ensures that the model can recognize the relative positions along a single axis and the crucial interactions between patches across both dimensions.
+We show our 2D form in Figure 2 (d).
 In our implementation, we assign horizontal and vertical positions with independent dimensions to store the Gray Code.
 Formally, the attention map $\mathbf{AttnMap}$ is:
-$\small\begin{split}\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot(&[\mathbf{Q}\parallel G(\mathbf{h})\parallel G(\mathbf{w})]\\ \oplus&[\mathbf{K}\parallel G(\mathbf{h})\parallel G(\mathbf{w})]).\end{split}$ (16)
-Here, $\mathbf{h}$ is the array of position indexes, specifically $\mathbf{h}=[0,1,2,\dots,h-1]$, where $h$ denotes the maximum patch index along the height axis.
+$\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot\left([\mathbf{Q}\parallel G(\mathbf{h})\parallel G(\mathbf{w})]\oplus[\mathbf{K}\parallel G(\mathbf{h})\parallel G(\mathbf{w})]\right).$ (12)
+Here, $\mathbf{h}$ is the array of position indices, specifically $\mathbf{h}=[0,1,2,\dots,h-1]$, where $h$ denotes the maximum patch index along the height axis.
 Similarly, $\mathbf{w}$ is along the width axis.
-We show our 2D form in Figure 2 (d).
+As for the 2D form of Log-PE, we can add $\mathbf{R^{h}}_{i,j}$ and $\mathbf{R^{w}}_{i,j}$ on $\mathbf{AttnMap}$, replacing the sequence length $L$ in Equation 11 with $h$ or $w$.
+However, in our pre-experiments, we found that spiking Transformers with Log-2D failed to converge due to the excessive magnitude.
+Therefore, we abandon the 2D form of Log-PE.
+5 Experiments
+5.1 Datasets
+To evaluate the RPE capabilities of the compared models, we conduct experiments on two sequential tasks: time-series forecasting and text classification.
+Following [6], we choose $4$ real-world datasets for time-series forecasting: Metr-la [27], Pems-bay [27], Electricity [28], Solar [28].
+For text classification, we follow [7] and conduct experiments on six benchmark datasets: Movie Reviews [29], SST-2 [30], SST-5, Subj, ChnSenti, and Waimai.
+Additionally, to demonstrate the versatility of our RPE method in image processing, we perform patch-based image classification experiments on two static datasets, CIFAR and Tiny-ImageNet, and one neuromorphic dataset, CIFAR10-DVS [2].
+The details of these datasets, metrics, and training hyperparameters are provided in Appendix D.
 Table 1: Experimental results of time-series forecasting on $4$ benchmarks with various prediction lengths $6,24,48,96$.
@@
 “R” denotes relative PE, while “A” denotes absolute PE.
-“w/o” denotes “without” while “w/” denotes “with”.
-The best results for each series of SNNs are highlighted in bold font.
-$\uparrow$ (indicating "higher is better") and $\downarrow$ (indicating "lower is better") are used to denote performance metrics.
+“w/” denotes “with”.
+The best results for each series of spiking Transformers are highlighted in bold font.
+$\uparrow$ ($\downarrow$) indicates that the higher (lower) the better.
 Results highlighted with shading are ours.
 All results are averaged across $3$ random seeds.
-Models PE Metric Metr-la (Short) Pems-bay (Short) Solar (Long) Electricity (Long) Avg.
+Models PE Metric Metr-la ($L=12$) Pems-bay ($L=12$) Solar ($L=168$) Electricity ($L=168$) Avg.
 Spike Type $6$ $24$ $48$ $96$ $6$ $24$ $48$ $96$ $6$ $24$ $48$ $96$ $6$ $24$ $48$ $96$
-Transformer w/ RPE R R2$\uparrow$ .$\bf.729$$\bf.560$$\bf.416$$\bf.306$$\bf.787$$730$$\bf.694$$\bf.676$$951$$854$$\bf.763$$\bf.720$$\bf.984$$\bf.978$$\bf.974$$\bf.968$$\bf.756$
+Transformer w/ RoPE R R2$\uparrow$ .$\bf.729$$\bf.560$$\bf.416$$\bf.306$$\bf.787$$730$$\bf.694$$\bf.676$$951$$854$$\bf.763$$\bf.720$$\bf.984$$\bf.978$$\bf.974$$\bf.968$$\bf.756$
 […]
@@
 […]
-Spikformer w/o PE R2$\uparrow$ .$697$$491$$383$$242$$768$$684$$678$$663$$903$$819$$715$$656$$956$$955$$953$$943$719.719.719.719
+Spikformer w/ Conv-PE (Original) A R2$\uparrow$ .$713$$527$$399$$267$$773$$697$$686$$667$$929$$828$$744$$674$$959$955.955.955.955 .955.955.955.955 .$954$733.733.733.733
 […]
-Spikformer w/ Conv-PE (Original) A R2$\uparrow$ .$713$$527$$399$$267$$773$$697$$686$$667$$929$$828$$744$$674$$959$955.955.955.955 .955.955.955.955 .$954$733.733.733.733
+Spikformer w/ RoPE R2$\uparrow$ .$699$$493$$390$$243$$768$$699$$680$$664$$911$$820$$714$$644$$954$$951$$949$$940$720.720.720.720
 […]
@@
 […]
-Spikformer-XNOR w/o PE R2$\uparrow$ .$699$$488$$380$$244$$764$$680$$679$$666$$906$$820$$721$$653$$960$$958$$951$$943$720.720.720.720
-[…]
 Spikformer-XNOR w/ Conv-PE A R2$\uparrow$ .$718$$531$$405$$269$$771$$693$$690$$665$$928$$829$$740$$669$$960$$957$$955$$953$733.733.733.733
-[…]
-Spikformer-XNOR w/ Binary-PE R2$\uparrow$ .$701$$488$$384$$252$$759$$698$$675$$644$$918$$815$$707$$677$$964$$959$$949$$945$721.721.721.721
 […]
@@
 […]
-QKFormer w/o PE (Original) R2$\uparrow$ .$702$$501$$377$$240$$750$$702$$664$$655$$905$$751$$497$$375$$964$$963$$958$$955$685.685.685.685
+SDT-V1 w/ Conv-PE (Original) R2$\uparrow$ .$689$$517$$409$$253$$769$$700$$647$$630$$917$$819$$723$$655$$956$$952$$949$$950$721.721.721.721
 […]
-QKFormer w/ Conv-PE A R2$\uparrow$ .$717$$513$$376$$246$$767$$706$$681$$654$$920$$748$$512$$416$$970$$967$$963$$958$695.695.695.695
+SDT-V1 w/ CPG-PE R2$\uparrow$ .$701$$525$$\bf.418$$257$$778$$\bf.716$$660$$\bf.656$$919$$\bf.820$$710$$644$$963$$960$$958$$952$727.727.727.727
+[…]
+R2$\uparrow$ .$\bf.714$$\bf.531$$415$$\bf.265$$\bf.784$$709$$\bf.672$$654$$\bf.921$$\bf.820$$\bf.730$$\bf.674$$\bf.972$$\bf.968$$\bf.963$$\bf.957$$\bf.734$
+[…]
+QKFormer w/ Conv-PE (Original) A R2$\uparrow$ .$717$$513$$376$$246$$767$$706$$681$$654$$920$$748$$512$$416$$970$$967$$963$$958$695.695.695.695
 […]
@@
 […]
-5 Experiments
-5.1 Datasets
-To evaluate the RPE capabilities of the compared models, we conduct experiments on two sequential tasks: time-series forecasting and text classification.
-Following Lv et al.
-(2024b), we choose $4$ real-world datasets for time-series forecasting: Metr-la (Li et al., 2017b), Pems-bay (Li et al., 2017b), Electricity (Lai et al., 2018), Solar (Lai et al., 2018).
-For text classification, we follow Lv et al.
-(2024a) and conduct experiments on six benchmark datasets: Movie Reviews (Pang & Lee, 2005), SST-2 (Socher et al., 2013), SST-5, Subj, ChnSenti, and Waimai.
-Additionally, to demonstrate the versatility of our RPE method in image processing, we perform patch-based image classification experiments on one static dataset, CIFAR, and one neuromorphic dataset, CIFAR10-DVS (Li et al., 2017a).
-The details of these datasets, metrics, and training hyper-parameters are provided in Appendix B.
 5.2 Time-Series Forecasting
-Time-series forecasting is a suitable task for evaluating the effectiveness of our proposed RPE method because it involves capturing the temporal dependencies between data points over time.
-We follow the framework outlined by Lv et al.
-(2024b) to conduct time-series forecasting experiments using SNN models.
-Specifically, we use Spikformer (Zhou et al., 2023b) and the current state-of-the-art (SOTA) model, QKFormer (Zhou et al., 2024), as the backbone architectures, both of which incorporate spiking self-attention mechanisms.
-We modify the spiking self-attention mechanism as outlined in Section 4.2 to create two variants: Spikformer-XNOR and QKFormer-XNOR.
-Subsequently, we integrate our proposed Gray-PE and its extended forms into these SNN architectures.
-For the baseline comparisons, we consider the following: (1) spiking absolute PE and CPG-PE (Lv et al., 2024a), (2) the PE used in the original architectures, referred to as Conv-PE, and (3) the integration of Binary-PE into the XNOR-based SSA.
-To isolate the effects of the XNOR modification, we also perform experiments on Spikformer-XNOR and QKFormer-XNOR without any positional encoding, as well as with Conv-PE.
+We follow the SeqSNN [6] framework to conduct time-series forecasting experiments.
+Specifically, we take Spikformer [4], Spike-driven Transformer (SDT) V1 [5], and the current visual state-of-the-art (SOTA) model, QKFormer [20], as the backbone architectures.
+We modify the SSA mechanism as outlined in Section 4.2 to create two variants: Spikformer-XNOR and QKFormer-XNOR.
+SDT adopts a variant of SSA, which makes it only able to integrate Log-PE but not Gray-PE.
 We present the performance of the compared SNN models with various positional encoding methods in Table 1.
 The key findings are as follows:
-Table 2: Accuracy (%) on $6$ text classification benchmarks.
-The best results of SNNs and ANNs are formatted in bold font.
-We set the sentence length to $128$ for the English dataset and $32$ for the Chinese dataset.
-Experimental results are averaged across $5$ random seeds.
+(1) Directly applying RPE methods to spiking Transformers is ineffective.
+Specifically, Spikformer with rotary position embedding (RoPE) exhibits poor performance across all benchmarks.
+As discussed in Section 1, we argue that this limitation stems from the binary nature of spiking neurons during the computation of $\mathbf{Q}$ and $\mathbf{K}$, which makes it difficult to disentangle positional information from sparse spiking activations.
+(2) The XNOR modification does not impact the performance of the original SNN models.
+The average performance of Spikformer with Conv-PE is nearly identical to that of Spikformer-XNOR with Conv-PE.
+This suggests that our XNOR modification of the SSA does not affect the performance of the original SNN models.
+(3) Gray-PE and Log-PE, enable spiking Transformers to achieve the best performance among their variants.
+CPG-PE is a spiking version of absolute PE designed for SNNs.
+Spikformer and QKFormer, when equipped with our proposed Gray-PE and Log-PE, consistently outperform all other corresponding variants.
+(4) For long input sequences, Log-PE is more effective than Gray-PE in capturing relative positional information.
+The input sequence length for Metr-la and Pems-bay is $12$, whereas for Solar and Electricity, it is $168$.
+On the long-sequence datasets Solar and Electricity, spiking Transformers equipped with Log-PE consistently outperform those with Gray-PE across nearly all prediction length settings.
+This result indicates that Log-PE is more effective for processing long input sequences.
+5.3 Text Classification
+We conduct experiments to assess the efficacy of spiking Transformers with Gray-PE and Log-PE in text classification tasks.
+By comparing them against alternative PE techniques, we demonstrate their superior ability to model complex linguistic structures and contextual dependencies.
+Our experimental setup strictly adheres to the methodology outlined in [7], and the results are shown in Table 2.
+Table 2: Accuracy (%) on $6$ text classification benchmarks. Note that QKFormers fail to converge in the text classification task. Experimental results are averaged across $5$ random seeds.
 Model PE Param(M) English Dataset (Length = $128$) Chinese Dataset (Length = $32$) Avg.
@@
 Spikformer-XNOR w/ Log-PE R $109.8$ $\bf 83.88$±$\pm 0.40$ $\bf 84.64$±$\pm 0.37$ $\bf 92.80$±$\pm 0.30$ $\bf 44.52$±$\pm 0.43$ $\bf 87.95$±$\pm 0.34$ $88.46$±$\pm 0.28$ $\bf 80.38$
-(1) Relative PE generally outperforms absolute PE in time-series forecasting tasks.
-Based on the comparison between the Transformer with RPE and sinusoidal-PE, as well as Spikformer-XNOR with Conv-PE and Gray-PE, we observe that relative PE consistently outperforms absolute PE, particularly on long-term time-series datasets such as Solar and Electricity.
-(2) The XNOR modification does not impact the performance of the original SNN models.
-The average performance of Spikformer without PE is nearly identical to that of Spikformer-XNOR without PE.
-This suggests that our XNOR modification of the SSA does not affect the performance of the original SNN models.
-Consequently, we can attribute the observed performance improvements solely to the introduced positional encoding techniques, excluding any impact from the XNOR operation itself.
-(3) Gray-PE and its extended form, Log-PE, enable Spikformer and QKFormer to achieve the best performance among their variants.
-CPG-PE is a spiking version of absolute PE designed for SNNs.
-Spikformer and QKFormer, when equipped with our proposed Gray-PE and Log-PE, consistently outperform all other corresponding variants, including CPG-PE and Binary-PE.
-Furthermore, we were surprised to find that Binary-PE fails to capture the relative positional information.
-This leads us to reconsider the underlying issue: if a binary encoding method cannot represent distances between certain positions, it is unsuitable as an RPE solution for SNNs.
-In contrast, Gray Code and Log-PE provide approximate binary encoding approaches that “partially” achieve this goal.
-(4) For long input sequences, Log-PE is more effective than Gray-PE in capturing relative positional information.
-The input sequence length for Metr-la and Pems-bay is $12$, whereas, for Solar and Electricity, it is $168$.
-As an approximated approach of RPE, Gray-PE can only represent the distance between two indexes differing by $2^{n}$.
-This limitation means that SNNs using Gray-PE may experience a performance drop when processing longer input sequences.
-Furthermore, on the long-sequence datasets Solar and Electricity, SNNs equipped with Log-PE consistently outperform those with Gray-PE across nearly all prediction length settings.
-For instance, on the Solar dataset with a prediction length of $48$, QKFormer with Log-PE significantly outperforms QKFormer with Gray-PE, showing an average improvement of $0.038$ in R2.
-This result indicates that Log-PE is more effective for processing long input sequences.
-5.3 Text Classification
-Text classification is another representative task for which RPE plays an important role.
-Therefore, we conduct experiments to evaluate the performance of SNNs with RPE in text classification tasks, comparing it against other PE techniques to highlight its advantages in capturing complex linguistic structures and contextual nuances.
-For this part, we strictly follow the experimental setup used in the text classification experiments in Lv et al.
-(2024a), and the results are presented in Table 2.
 Based on the results in Table 2, it is evident that our proposed Gray-PE and Log-PE significantly outperform the other spiking positional encoding methods across several key benchmarks.
@@
 Unlike spiking absolute PE, i.e., CPG-PE, which struggles to adapt to the nuances of language, Gray-PE and Log-PE provide a more flexible and context-sensitive representation, improving the model’s ability to classify sentences accurately.
+5.4 Patch-based Image Classification
 Table 3: Accuracy (%) on image classification benchmarks.
-Numbers with denote our implementation, representing results that differ from those reported in the original papers.
+Numbers with denote our implementations.
 The best and second-best results are highlighted in bold and underlined formats, respectively.
-“R” denotes relative PE, while “A” denotes absolute PE.
 The results with shading are ours.
 Results are averaged across $4$ random seeds.
-Model PE CIFAR10 CIFAR10-DVS CIFAR100 Avg.
-Spike Type Param (M) Accuracy Param (M) Accuracy Param (M) Accuracy
-Vision-Transformer A $9.32$ $\bm{96.73}$ $9.36$ $\bm{81.02}$
-Spikformer w/o PE $8.00$ $93.98$ $1.99$ $76.60$ $8.04$ $74.02$ $81.53$
-Spikformer w/ Conv-PE (Original) A $9.32$   $\;\;94.80^{*}$ $2.57$   $\;\;78.10^{*}$ $9.36$   $\;\;77.04^{*}$ $83.31$
-Spikformer w/ CPG-PE A $8.17$ $95.06$ $2.06$ $\underline{78.40}$ $8.20$ $77.82$ $83.76$
-Spikformer-XNOR w/o PE $8.00$ $94.14$ $1.99$ $76.70$ $8.04$ $74.16$ $81.67$
-Spikformer-XNOR w/ Conv-PE A $9.32$ $95.20$ $2.57$ $77.90$ $9.36$ $78.04$ $83.71$
-Spikformer-XNOR w/ Gray-PE 1D R $8.00$ $\underline{95.46}$ $1.99$ $77.90$ $8.04$ $\underline{78.12}$ $\underline{83.83}$
-Spikformer-XNOR w/ Gray-PE 2D R $8.00$ $\bf 95.66$ $1.99$ $\bf 78.70$ $8.04$ $\bf 78.45$ $\bf 84.27$
-QKFormer w/o PE (Original) $6.74$ $\;\;\underline{96.32}^{*}$ $1.50$   $\;\;\bf 83.40^{*}$ $6.76$   $\;\;\bf 80.90^{*}$ $\bf 86.87$
-QKFormer w/ CPG-PE A $7.01$ $96.30$ $1.58$ $82.00$ $7.04$ $80.52$ $86.27$
-QKFormer-XNOR w/o PE $6.74$ $96.28$ $1.50$ $82.90$ $6.76$ $80.80$ $86.66$
-QKFormer-XNOR w/ Gray-PE 1D R $6.74$ $96.22$ $1.50$ $82.20$ $6.76$ $80.48$ $86.30$
-QKFormer-XNOR w/ Gray-PE 2D R $6.74$ $\bf 96.36$ $1.50$ $\underline{83.10}$ $6.76$ $\underline{80.82}$ $\underline{86.76}$
-5.4 Patch-based Image Classification
-We also conduct experiments on ViT-based SNNs, including Spikformer and the current SOTA model, QKFormer.
-As discussed in Section 4.5, we extend Gray-PE into a two-dimensional form and integrate it into the patch-based framework to enhance the spatial relationship understanding of spiking Transformers.
-The results of these experiments are presented in Table 3, where we draw conclusions that:
-(1) Gray-PE enhances the performance of Spikformer with less parameters.
-Gray-PE, whether in 1D or 2D form, consistently improves Spikformer’s classification accuracy.
-As an approximation of RPE, Gray-PE successfully outperforms spiking absolute PE (i.e., CPG-PE), highlighting the strong dependency between patches within an image.
-Notably, Spikformer-XNOR with Gray-PE 2D achieves the highest performance across all three datasets, demonstrating the effectiveness of 2D form Gray-PE in capturing relative positional information.
-(2) On image classification, QKFormer is insensitive to positional encoding.
-Surprisingly, QKFormer shows limited improvements or even performance drops when using PE techniques.
-Whether with CPG-PE or Gray-PE, QKFormer fails to achieve significant performance gains and, in fact, shows a decline in performance.
-We attribute this phenomenon to the special attention mechanism in QKFormer, known as “q-k attention”, which sums the query matrix along the channel dimension before performing a dot product with key matrix.
-This strategy eliminates the positional representation provided by both CPG-PE and Gray-PE, which negatively impacts performance.
-(3) For Gray-PE, the 2D form performs better than the 1D form on processing image patches.
-Through a comparison of the performance of SNNs using Gray-PE 1D and 2D, we find that the design of the two-dimensional form of Gray Code is effective.
-In the case of Spikformers, Gray-PE 2D significantly outperforms Gray-PE 1D, with a $0.44\%$ accuracy improvement averagely.
-Even for QKFormer, where PE’s influence is limited, Gray-PE 2D still far outperforms its 1D variant.
-These results confirm that Gray-PE 2D is superior to Gray-PE 1D in handling positional information for image classification tasks, validating the success of our proposed Gray-PE 2D.
-5.5 Positional Encoding Analysis
-In this section, we will analyze the following points: (1) Impact of the number of bits on the performance of SNNs when using Gray-PE.
-(2) Rationale behind why we use an approximate RPE rather than a complete one in SNNs.
+Model PE CIFAR10 CIFAR10-DVS CIFAR100 Tiny-ImageNet Avg.
+Spike Type Param (M) Acc Param (M) Acc Param (M) Acc Param (M) Acc
+Vision-Transformer A $9.32$ $\bm{96.73}$ $9.36$ $\bm{81.02}$ $9.40$ $\bm{62.18}$
+Spikformer w/ Conv-PE (Original) A $9.32$   $\;\;94.80^{*}$ $2.57$   $\;\;78.10^{*}$ $9.36$   $\;\;77.04^{*}$ $9.40$   $\;\;48.10^{*}$ $74.51$
+Spikformer w/ CPG-PE A $8.17$ $95.06$ $2.06$ $\underline{78.40}$ $8.20$ $77.82$ $8.24$ $\;\;\underline{48.52}^{*}$ $\underline{74.95}$
+Spikformer-XNOR w/ Gray-PE 1D R $8.00$ $\underline{95.46}$ $1.99$ $77.90$ $8.04$ $\underline{78.12}$ $8.08$ $48.33$ $\underline{74.95}$
+Spikformer-XNOR w/ Gray-PE 2D R $8.00$ $\bf 95.66$ $1.99$ $\bf 78.70$ $8.04$ $\bf 78.45$ $8.08$ $\bf 48.74$ $\bf 75.39$
+In this section, we evaluate ViT-based SNNs, Spikformer, which adopts a patch-splitting processing approach.
+To enhance compatibility with this framework, we extend Gray-PE into a 2D form and integrate it into the patch-based architecture.
+The experimental results are summarized in Table 3.
+We draw conclusions that:
+(1) Gray-PE enhances the performance of Spikformer while maintaining parameter efficiency.
+Both 1D and 2D variants of Gray-PE consistently improve classification accuracy.
+Notably, Gray-PE surpasses spiking absolute PE (CPG-PE), indicating its superior ability to model inter-patch dependencies within images, even as an approximation of RPE.
+(2) The 2D variant of Gray-PE demonstrates superior performance over its 1D counterpart in processing image patches.
+Empirical comparisons between Spikformers equipped with Gray-PE 1D and 2D reveal that the two-dimensional form is highly effective.
+Specifically, Gray-PE 2D achieves an average accuracy improvement of $0.44\%$ over Gray-PE 1D.
+Furthermore, we present the image classification performance of the state-of-the-art QKFormer integrated with our proposed RPE methods in Appendix C.
+5.5 Capability of Processing Long Sequences
+In this section, we assess the effectiveness of our proposed relative positional encoding methods in handling long sequences within spiking Transformers.
+To this end, we use two text classification datasets characterized by long input samples: AGNEWS [31] and IMDB [32].
+Following [33], we fix the sequence max length to $1024$ for AGNEWS and $2048$ for IMDB.
+We train the Spikformer model using various positional encoding strategies on these datasets, and present the results in Table 4.
+Table 4: Accuracy (%) on $2$ long text classification benchmarks. We set the sentence length to $1024$ for AGNEWS and $2048$ for IMDB.
+Model PE AGNEWS IMDB Avg.
+Spike Type
+Fine-tuned BERT A $\bf 94.50$ $\bf 92.10$ $\bf 93.30$
+Spikformer w/ Conv-PE (Original) A $83.84$ $79.08$ $81.46$
+Spikformer w/ CPG-PE A $84.70$ $79.47$ $82.09$
+Spikformer-XNOR w/ Gray-PE R $84.92$ $79.79$ $82.36$
+Spikformer-XNOR w/ Log-PE R $\bf 86.77$ $\bf 80.46$ $\bf 83.62$
+As shown in Table 4, although Spikformer models lag behind the fine-tuned BERT in overall performance, both Log-PE and Gray-PE demonstrate effectiveness when handling long input sequences.
+Notably, Log-PE yields substantial performance improvements, suggesting its strong suitability for processing long texts.
+This outcome is expected, as Log-PE is specifically designed to accommodate long-range dependencies.
+5.6 Analysis and Ablation
+In this section, we analyze the following aspects: (1) The influence of internal properties in Gray-PE, (2) Ablation studies on XNOR and Log-PE (shown in Appendix B).
 Figure 3: Spikformer-XNOR with Gray-PE across various bit numbers ranging from $2$ to $12$ on (a) time-series forecasting tasks and (b) text classification tasks.
-For the first question, we present the results of our parameter sensitivity experiments in Figure 3.
-Consider that: if the number of bits used for encoding relative positions in Gray-PE is $b$, then the total number of unique encodings possible is $2^{b}$.
-We set the maximum sequence length is $L$, so relative distances range from $1$ to $L-1$.
-According to the pigeonhole principle, if $L-1>2^{b}$, there will inevitably be at least two distances that are represented identically.
+Consider that: If the number of bits used for encoding relative positions in Gray Code is $b$, then the total number of unique encodings possible is $2^{b}$.
+We set the maximum sequence length is $L$, so relative distances range from 00 to $L-1$.
+According to the pigeonhole principle, if $L-1>2^{b}$, there will be at least two distances that are represented identically.
 This issue can be mitigated by increasing $b$ to cover the range of relative distances up to $L-1$.
@@
 However, for shorter datasets like Metr-la (Length $=12$) and ChnSenti (Length $=32$), $5$ bits are sufficient.
-Table 4: Time-series forecasting performance of SNNs trained with Complete RPE (C-RPE) method. We take Metr-la and Electricity as examples. denotes failure to converge.
+6 Conclusion
+In this work, we have designed several RPE methods for spike Transformers.
+Our approach preserves the spiking nature of SNNs while effectively representing relative positions.
+Experimental evaluations on time series forecasting, text classification, and image classification demonstrate significant performance improvements.
+These empirical results, together with theoretical analysis of the proposed RPE methods, highlight the potential to enhance the versatility and applicability of SNNs across various domains.
+Future work and limitations are discussed in Appendix E.
+Appendix A Proof of Theorem 1
+This section provides a detailed mathematical proof of Theorem 1. We use the standard Reflected Binary Code (RBC) $G(x)$ as our Gray Code:
+Definition 1.
+The Reflected Binary Code (Gray Code) of an integer $x$ is defined as:
+$G(x)=x\oplus(x\gg 1),$ (13)
+where $\oplus$ denotes the bitwise XOR operation, and $\gg$ denotes the arithmetic right shift.
+and we restate Theorem 1 here:
+Theorem 1.
+For any non-negative integer $n$, and for any pair of decimal integers $a$ and $b=a+2^{n}$, the Hamming distance between their Gray Code representations $G(a)$ and $G(b)$ is consistently:
+$d_{H}(G(a),G(b))=\begin{cases}1&\text{if }n=0,\\ 2&\text{if }n\geq 1.\end{cases}$ (14)
+Here, Hamming distance is the number of different bits between two binary representations.
+Proof.
+For $n\geq 1$, consider $b=a+2^{n}$. We analyze the XOR of their Gray Codes:
+$G(a)\oplus G(a+2^{n})=\left[a\oplus\left(a\gg 1\right)\right]\oplus\left[(a+2^{n})\oplus\left((a+2^{n})\gg 1\right)\right].$ (15)
+Using the associativity and commutativity of XOR, we regroup:
+$G(a)\oplus G(a+2^{n})=\left[a\oplus\left(a+2^{n}\right)\right]\oplus\left[(a\gg 1)\oplus\left((a+2^{n})\gg 1\right)\right].$ (16)
+Let us denote:
+$\Delta_{1}=a\oplus\left(a+2^{n}\right),\quad\Delta_{2}=(a\gg 1)\oplus\left(\left(a+2^{n}\right)\gg 1\right).$ (17)
+Case 1: $n=0$
+In this case, $b=a+1$, and it is well known that adjacent integers in the Gray code differ by exactly one bit. Therefore, we have $d_{H}(G(a),G(b))=d_{H}(G(a),G(a+1))=1$.
+Case 2: $n\geq 1$
+We consider two subcases based on the bit at position $n$ in $a$.
+Subcase A: Bit $n$ in $a$ is 0
+Then $a+2^{n}$ flips bit $n$ from 0 to 1, with no carry. Hence:
+$\Delta_{1}=2^{n},\quad\Delta_{2}=2^{n-1}.$ (18)
+Therefore,
+$G(a)\oplus G(b)=2^{n}\oplus 2^{n-1}.$ (19)
+This value has exactly two bits set (at positions $n$ and $n-1$), so the Hamming distance is 2.
+Subcase B: Bit $n$ in $a$ is 1
+Then adding $2^{n}$ to $a$ causes a carry from bit $n$ upwards. Let $c$ be the smallest index greater than $n$ such that bit $c$ in $a$ is 00; bits $n$ through $c-1$ are all $1$. Then:
+$\Delta_{1}=a\oplus(a+2^{n})=\sum_{i=n}^{c}2^{i}=\underbrace{1\dots 1}_{c-n+1}\underbrace{0\dots 0}_{n}\ {}_{(2)},$ (20)
+has ones at bits $n$ through $c$. We denote $\cdot_{(2)}$ for binary representation. Similarly,
+$\Delta_{2}=(a\gg 1)\oplus((a+2^{n})\gg 1)=\sum_{i=n-1}^{c-1}2^{i}=\underbrace{1\dots 1}_{c-n+1}\underbrace{0\dots 0}_{n-1}\ {}_{(2)},$ (21)
+has ones at bits $n-1$ through $c-1$. Thus,
+$G(a)\oplus G(b)=\Delta_{1}\oplus\Delta_{2}=(\sum_{i=n}^{c}2^{i})\oplus(\sum_{i=n-1}^{c-1}2^{i})=2^{c}+2^{n}$ (22)
+has ones at only positions $c$ and $n-1$, and all other bits are canceled due to alignment. The result has exactly two bits set, so the Hamming distance is $2$.
+Combining all cases, we conclude that for any non-negative integer $n$:
+$d_{H}(G(a),G(a+2^{n}))=\begin{cases}1&\text{if }n=0,\\ 2&\text{if }n\geq 1.\end{cases}$ (23)
+This rigorously proves the observed property of Gray Codes concerning the Hamming distance between numbers differing by powers of two.
+Appendix B Ablation Study on Log-PE and XNOR
+In this section, we compare the performance of vanilla spiking Transformers with Log-PE, XNOR variants with Log-PE, and models equipped with complete relative positional encoding (C-RPE).
+As discussed in Section 4.4, C-RPE is implemented by setting $R_{i,j}=\frac{L-1}{|i-j|+1}$ and adding it directly to the attention scores.
+Table S1: Ablation study on XNOR and Log-PE.
+We take the time-series forecasting performance of SNNs on Metr-la and Electricity as examples.
+C-RPE denotes Complete RPE.
+$\uparrow$ ($\downarrow$) indicates that the higher (lower) the better.
+denotes failure to converge.
 Model (Prediction Length $=24$) Metr-la ($L=12$) Electricity ($L=168$)
 R2 $\uparrow$ RSE $\downarrow$ R2 $\uparrow$ RSE $\downarrow$
-Spikformer-XNOR w/o PE .$491$$753$$955$375.375.375.375
+Spikformer w/ Log-PE .$484$$763$$710^{*}.$ $1.03^{*}$
 Spikformer-XNOR w/ Log-PE .$\bf.535$$\bf.719$$\bf.974$$\bf.300$
 Spikformer-XNOR w/ C-RPE .$158^{*}.$$967^{*}.$$710^{*}.$ $1.03^{*}$
-QKFormer-XNOR w/o PE .$501$$740$$963$352.352.352.352
+QKFormer w/ Log-PE .$475$$788$$710^{*}.$ $1.03^{*}$
 QKFormer-XNOR w/ Log-PE .$\bf.541$$\bf.715$$\bf.974$$\bf.285$
 QKFormer-XNOR w/ C-RPE .$430$$824$$710^{*}.$ $1.03^{*}$
-Regarding the second question, we show the performance of SNNs using complete RPE (C-RPE).
-As discussed in Section 4.4, C-RPE is obtained by setting $R_{i,j}$ as $\frac{L-1}{|i-j|}$ and adding it to the attention map.
-However, the large values of this method can cause the vanilla attention map to fail in recording accurate similarity scores.
-As shown in Table 4, we observe that SNNs with C-RPE significantly underperform compared to other variants, with some even failing to converge, due to the detrimental impact of the large positional encoding values on the training process.
-Besides, we design another extended form of Gray-PE, leveraging hierarchical relative positional representation.
-However, we find that when an encoding method involves carry-over, ensuring the uniqueness of the spike representation for each relative distance becomes challenging.
-A detailed analysis of this issue can be found in Appendix C.
-6 Conclusion
-In this paper, we propose a series of relative positional encoding methods for spiking Transformers.
-Our approach preserves the spiking nature of SNNs while effectively representing relative distances.
-Experimental evaluations on time series forecasting, text classification, and image classification demonstrate significant performance improvements.
-These results validate the effectiveness of the proposed RPE method and highlight its potential to enhance the versatility and applicability of SNNs across various domains.
-Future work and limitations are discussed in Appendix D.
-Impact Statement
-This research enhances spiking neural networks by introducing an approximation method for relative positional encoding method based on Gray Code.
-The proposed approach broadens the applicability of SNNs across diverse domains, enabling more robust and versatile neural architectures.
-Lastly, we do not think our work will negatively impact ethical aspects or future societal consequences.
-Appendix A Proof of Theorem 4.1
-This section provides a detailed mathematical proof of Theorem 4.1.
-A.1 Statement
-For any non-negative integer $n$, and for any pair of integers $a$ and $b=a+2^{n}$, the Hamming distance between their Gray Code representations $G(a)$ and $G(b)$ is:
-$d_{H}(G(a),G(b))=\begin{cases}1&\text{if }n=0,\\ 2&\text{if }n\geq 1.\end{cases}$ (17)
-A.2 Proof
-We aim to determine the Hamming distance $d_{H}(G(a),G(a+2^{n}))$.
-A.2.1 Case 1: $n=0$
-When $n=0$, $b=a+1$. By the definition of the Gray Code:
-$G(a)=a\oplus\left(a\gg 1\right),\quad G(a+1)=(a+1)\oplus\left((a+1)\gg 1\right).$ (18)
-The Gray Code inherently ensures that consecutive integers differ by exactly one bit. Therefore:
-$d_{H}(G(a),G(a+1))=1.$ (19)
-This establishes the base case for $n=0$.
-A.2.2 Case 2: $n\geq 1$
-For $n\geq 1$, consider $b=a+2^{n}$. We analyze the XOR of their Gray Codes:
-$\begin{split}&G(a)\oplus G(a+2^{n})\\ =&\left(a\oplus\left(a\gg 1\right)\right)\oplus\left((a+2^{n})\oplus\left((a+2^{n})\gg 1\right)\right).\end{split}$ (20)
-Rearranging the terms:
-$\begin{split}&G(a)\oplus G(a+2^{n})\\ =&\left(a\oplus(a+2^{n})\right)\oplus\left((a\gg 1)\oplus\left((a+2^{n})\gg 1\right)\right).\end{split}$ (21)
-Analyzing $a\oplus(a+2^{n})$: Adding $2^{n}$ to $a$ flips the $n$-th bit of $a$. Therefore:
-$a\oplus(a+2^{n})=2^{n}.$ (22)
-Analyzing $(a\gg 1)\oplus((a+2^{n})\gg 1)$: Shifting both $a$ and $a+2^{n}$ to the right by one bit:
-$(a+2^{n})\gg 1=\left(a\gg 1\right)+2^{n-1}.$ (23)
-Thus:
-$\begin{split}&(a\gg 1)\oplus\left((a+2^{n})\gg 1\right)\\ =&\left(a\gg 1\right)\oplus\left(\left(a\gg 1\right)+2^{n-1}\right)\\ =&2^{n-1}.\end{split}$ (24)
-Combining the Results:
-$G(a)\oplus G(a+2^{n})=2^{n}\oplus 2^{n-1}.$ (25)
-In binary representation, $2^{n}$ and $2^{n-1}$ have single 1’s in distinct positions. Their XOR results in a binary number with exactly two 1’s:
-$2^{n}\oplus 2^{n-1}=2^{n}+2^{n-1}.$ (26)
-Therefore, the Hamming weight (number of 1’s) of $G(a)\oplus G(a+2^{n})$ is $2$.
-Therefore, for $n\geq 1$, since the XOR of the Gray Codes yields a binary number with exactly two 1’s, the Hamming distance between $G(a)$ and $G(a+2^{n})$ is:
-$d_{H}(G(a),G(a+2^{n}))=2.$ (27)
-A.3 Final Conclusion
-Combining both cases, we conclude that for any non-negative integer $n$:
-$d_{H}(G(a),G(a+2^{n}))=\begin{cases}1&\text{if }n=0,\\ 2&\text{if }n\geq 1.\end{cases}$ (28)
-This rigorously proves the observed property of Gray Codes concerning the Hamming distance between numbers differing by powers of two.
-Appendix B Experimental Settings
-B.1 Datasets
-B.1.1 Time-series Forecasting
-We strictly follow the dataset settings of CPG-PE (Lv et al., 2024a).
-The datasets we used are as follows: Metr-la (Li et al., 2017b): Average traffic speed data collected from highways in Los Angeles County.
-Pems-bay (Li et al., 2017b): Average traffic speed data from the Bay Area.
-Electricity (Lai et al., 2018): Hourly electricity consumption data in kilowatt-hours (kWh) of $321$ clients.
-Solar (Lai et al., 2018): Solar power production.
+As shown in Table S1, both Spikformer and QKFormer with C-RPE perform significantly worse than their Log-PE counterparts, with some variants failing to converge entirely.
+This degradation is attributed to the overly large positional encodings disrupting the training dynamics.
+Furthermore, observed from vanilla Spikformer with Log-PE, we confirm that the dot product, which does not use Hamming distance to measure relative distance, is not suitable for RPE methods.
+Appendix C Performance of QKFormers on Image Classification
+In this section, we conduct experiments on current SOTA spiking Transformer, QKFormer [20].
+Table S2: Accuracy (%) of QKFormer on image classification benchmarks.
+Numbers with denote our implementations.
+“PE” stands for positional encoding.
+“R” denotes relative PE, while “A” denotes absolute PE.
+Results are averaged across $4$ random seeds.
+Model PE CIFAR10 CIFAR10-DVS CIFAR100 Tiny-ImageNet Avg.
+Spike Type Param (M) Acc Param (M) Acc Param (M) Acc Param (M) Acc
+Vision-Transformer A $9.32$ $\bm{96.73}$ $9.36$ $\bm{81.02}$ $9.40$ $\bm{62.18}$
+QKFormer w/ Conv PE (Original) $6.74$ $\;\;\underline{96.32}^{*}$ $1.50$   $\;\;\bf 83.40^{*}$ $6.76$   $\;\;\bf 80.90^{*}$ $6.78$   $\;\;\bf 58.07^{*}$ $\bf 79.67$
+QKFormer w/ CPG-PE A $7.01$ $96.30$ $1.58$ $82.00$ $7.04$ $80.52$ $7.08$   $\;\;56.75^{*}$ $78.89$
+QKFormer-XNOR w/ Gray-PE 1D R $6.02$ $96.22$ $1.41$ $82.20$ $6.04$ $80.48$ $6.06$ $57.21$ $79.03$
+QKFormer-XNOR w/ Gray-PE 2D R $6.02$ $\bf 96.36$ $1.41$ $\underline{83.10}$ $6.04$ $\underline{80.82}$ $6.06$ $\underline{57.94}$ $\underline{79.56}$
+As shown in Table S2, we find that: QKFormer exhibits insensitivity to positional encoding in image classification.
+QKFormer exhibits minimal performance gains, or even degradation, when augmented with PE techniques, including both CPG-PE and Gray-PE.
+We attribute this to its attention design, which aggregates queries along the channel dimension before the dot product with keys.
+This design inherently biases the model toward spatially specific features while suppressing temporal dependencies.
+Previous studies [34, 35, 36, 37] have shown that patch-based image classification primarily focuses on spatial (i.e., channel-wise) information rather than the sequential dependencies between patches.
+This contrasts with sequence modeling tasks such as time-series forecasting and text classification, where capturing inter-token dependencies is crucial.
+In image classification, our positional encoding encourages the model to emphasize sequential relationships between patches, which introduces a conflict with the QKFormer’s attention mechanism, ultimately hindering performance in this domain.
+Appendix D Experimental Settings
+D.1 Datasets
+D.1.1 Time-series Forecasting
+We strictly follow the dataset settings of CPG-PE [7].
+The datasets we used are as follows: Metr-la [27]: Average traffic speed data collected from highways in Los Angeles County.
+Pems-bay [27]: Average traffic speed data from the Bay Area.
+Electricity [28]: Hourly electricity consumption data in kilowatt-hours (kWh) of $321$ clients.
+Solar [28]: Solar power production.
 The detailed statistical characteristics and distribution ratios for each dataset are presented below:
-Table 5: The statistics of time-series datasets.
+Table S3: The statistics of time-series datasets.
 Dataset Samples Variables Observation Length Train-Valid-Test Ratio
@@
 Electricity $26,304$ $321$ $168,(\operatorname{long-term})$ $(0.6,0.2,0.2)$
-B.1.2 Text Classification
-For text classification, we follow Lv et al.
-(2023b) to conduct experiments on six easy discrimination tasks, covering both English and Chinese datasets.
+D.1.2 Text Classification
+For text classification, we follow [10] to conduct experiments on six easy discrimination tasks, covering both English and Chinese datasets.
 Here are the datasets we used in text classification experiments:
-The MR dataset, which stands for Movie Review, contains movie-review documents labeled based on their overall sentiment polarity (positive or negative) or subjective rating (Pang & Lee, 2005).
-SST-$5$ includes $11,855$ sentences from movie reviews for sentiment classification across five categories: very negative, negative, neutral, positive, and very positive (Socher et al., 2013).
+AGNEWS [31] is a large-scale text classification benchmark derived from AG’s corpus of news articles, containing $120,000$ training samples and $7,600$ valid samples evenly distributed across four categories—World, Sports, Business, and Science/Technology.
+IMDB [32] is a benchmark for binary sentiment classification, containing $50,000$ movie reviews labeled as positive or negative, split evenly into training and test sets to evaluate natural language understanding and opinion mining models.
+The MR dataset, which stands for Movie Review, contains movie-review documents labeled based on their overall sentiment polarity (positive or negative) or subjective rating [29].
+SST-$5$ includes $11,855$ sentences from movie reviews for sentiment classification across five categories: very negative, negative, neutral, positive, and very positive [30].
 SST-$2$ is the binary version of SST-$5$, containing only two classes: positive and negative.
-The Subj dataset is designed to classify sentences as either subjective or objective [^1] https://www.cs.cornell.edu/people/pabo/movie-review-data/.
-ChnSenti consists of approximately $7,000$ Chinese hotel reviews, each annotated with a positive or negative label [^2] https://raw.githubusercontent.com/SophonPlus/ChineseNlpCorpus/master/datasets/ChnSentiCorp_htl_all/ChnSentiCorp_htl_all.csv.
-Waimai contains around $12,000$ Chinese user reviews from a food delivery platform, intended for binary sentiment classification (positive and negative) [^3] https://raw.githubusercontent.com/SophonPlus/ChineseNlpCorpus/master/datasets/waimai_10k/waimai_10k.csv.
-B.1.3 Image Classification
+The Subj dataset is designed to classify sentences as either subjective or objective***https://www.cs.cornell.edu/people/pabo/movie-review-data/.
+ChnSenti consists of approximately $7,000$ Chinese hotel reviews, each annotated with a positive or negative labelhttps://raw.githubusercontent.com/SophonPlus/ChineseNlpCorpus/master/datasets/ChnSentiCorp_htl_all/ChnSentiCorp_htl_all.csv.
+Waimai contains around $12,000$ Chinese user reviews from a food delivery platform, intended for binary sentiment classification (positive and negative)https://raw.githubusercontent.com/SophonPlus/ChineseNlpCorpus/master/datasets/waimai_10k/waimai_10k.csv.
+D.1.3 Image Classification
 Here are the datasets we used in image classification experiments:
@@
 The relatively low resolution of the images makes the dataset a challenging benchmark for evaluating model performance in small-scale image classification tasks.
+The Tiny-ImageNet dataset is a simplified subset of the original ImageNet, designed for efficient experimentation in image classification and deep learning research.
+It consists of $200$ object classes, each containing $500$ training images, $50$ validation images, and $50$ test images (totaling $100,000$ training, $10,000$ validation, and $10,000$ test images).
+All images are downsampled to a resolution of $64\times 64$ pixels, balancing computational feasibility with visual complexity Designed for efficient deep learning research, it reduces computational costs while maintaining diversity.
 The CIFAR10-DVS dataset represents a neuromorphic adaptation of the original CIFAR10 set, where static images have been converted into dynamic representations that simulate the recording capabilities of a Dynamic Vision Sensor (DVS) camera.
@@
 This dataset is particularly valuable for testing models designed for neuromorphic systems and event-based vision tasks, offering a more realistic and biologically plausible approach to image classification.
-B.2 Time-series Forecasting
+D.2 Time-series Forecasting
 Metrices
@@
 Training Hyper-parameters
-we set the training batch size as $32$ and adopt Adam (Kingma & Ba, 2014) optimizer with a cosine scheduler of learning rate $1\times 10^{-4}$.
+we set the training batch size as $32$ and adopt Adam [38] optimizer with a cosine scheduler of learning rate $1\times 10^{-4}$.
 An early stopping strategy with a tolerance of $30$ epochs is adopted.
-For other configurations, we honestly follow the SeqSNN framework  [^4] https://github.com/microsoft/SeqSNN proposed by Lv et al.
-(2024b).
+For other configurations, we honestly follow the SeqSNN framework §§§https://github.com/microsoft/SeqSNN proposed by [6].
 We conducted time-series forecasting experiments on 24G-V100 GPUs.
 On average, a single experiment takes about $1$ hour under the settings above.
-B.3 Text Classification
+D.3 Text Classification
 Model Achirecture
 All Spikformers are with $12$ encoder blocks and $768$ feature embedding dimension.
-We have substituted layer normalization of SpikeBERT (Lv et al., 2023a) with batch normalization in our directly-trained Spikformer models for text classification tasks.
+We have substituted layer normalization of SpikeBERT [39] with batch normalization in our directly-trained Spikformer models for text classification tasks.
 Training Hyper-parameters
 We directly trained Spikformers with arctangent surrogate gradients on all datasets.
-We use the BERT-Tokenizer in Huggingface [^5] https://huggingface.co/ to tokenize the sentences to token sequences.
+We use the BERT-Tokenizer in Huggingfacehttps://huggingface.co/ to tokenize the sentences to token sequences.
 We pad all samples to the same sequence length of $256$.
-We conducted text classification experiments on $4$ RTX-3090 GPUs, and set the batch size as $32$, optimizer as AdamW (Loshchilov & Hutter, 2018) with weight decay of $5\times 10^{-3}$, and set a cosine scheduler of starting learning rate of $5\times 10^{-4}$.
+We conducted text classification experiments on $4$ RTX-3090 GPUs, and set the batch size as $32$, optimizer as AdamW [40] with weight decay of $5\times 10^{-3}$, and set a cosine scheduler of starting learning rate of $5\times 10^{-4}$.
 What’s more, in order to speed up the training stage, we adopt the automatic mixed precision training strategy.
 On average, a single experiment takes about $1.5$ hours under the settings above.
-B.4 Image Classification
+D.4 Image Classification
 Model Architecture
@@
 Training Hyper-parameters
-We honestly follow the experimental settings in Spikformer (Zhou et al., 2023b) and QKFormer (Zhou et al., 2024), whose source code and configuration files are available at https://github.com/ZK-Zhou/spikformer and https://github.com/zhouchenlin2096/QKFormer.
+We honestly follow the experimental settings in Spikformer [4] and QKFormer [20], whose source code and configuration files are available at https://github.com/ZK-Zhou/spikformer and https://github.com/zhouchenlin2096/QKFormer.
 As the training epochs are quite big ($300$ or $400$ epochs) in their settings, we choose to use one 80G-A100 GPU, and it takes about $3$ hours to conduct a single experiment, on average.
-Appendix C Analysis on Carry-Over
-We observe that when the encoding method involves carry-over, as in binary systems, it becomes difficult to ensure the uniqueness of the spike representation for each relative distance, especially with limited dimensions in the query and key vectors during self-attention.
-The carry-over operation introduces dependencies between bits, meaning the encoding of one bit can influence the encoding of higher-order bits.
-Consequently, the Hamming distance between the encodings of different relative distances is not fixed and can vary significantly due to carry-over.
-These dependencies also increase the likelihood of encoding collisions, where different relative distances might share identical or similar binary encodings.
-This can be empirically demonstrated by the poor performance of SNNs with Binary-PE in time-series forecasting, as shown in Table 1.
-From this perspective, Gray-PE and its extended variants can be viewed as approximate solutions, as they partially address the challenge of ensuring unique spike representations for a range of relative distances.
-Appendix D Limitations and Future Work
-D.1 Limitations
-Despite the promising enhancements introduced by our Gray Code-based relative positional encoding method for spiking neural networks, several limitations must be acknowledged.
-Firstly, the current Gray Code implementation may encounter scalability issues when applied to very long input sequences or more complex network architectures.
-Additionally, while Gray Code effectively preserves binary spike representations, it may limit the flexibility and adaptability of the encoding scheme across diverse data modalities and task requirements.
-Furthermore, our evaluations have been confined to specific applications such as time series forecasting, text classification, and image patch classification, which may not fully capture the method’s performance in other domains, such as object detection (Luo et al., 2024) and real-world geometry representation (Liao et al., 2024).
-D.2 Future Work
+Appendix E Limitations and Future Work
+E.1 Limitations
+Despite the promising enhancements introduced by our relative positional encoding method for spiking Transformers, several limitations must be acknowledged.
+Firstly, the current implementation may encounter scalability issues when applied to extremely long input (such as ultra-long texts with the length of $10240$) sequences.
+Additionally, while Gray-PE and Log-PE effectively preserve binary spike representations, they may limit the flexibility and adaptability of the encoding scheme across diverse data modalities and task requirements.
+Furthermore, our evaluations have been confined to specific applications such as time series forecasting, text classification, and image patch classification, which may not fully capture the method’s performance in other domains, such as object detection [41] and real-world geometry representation [42].
+E.2 Future Work
 Future work should focus on optimizing the Gray Code-based RPE to enhance its scalability and efficiency, enabling its deployment in larger and more intricate SNN models.
@@
 These efforts will contribute to the advancement of more versatile and powerful biologically inspired neural network architectures.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: We have clarified our claims in the abstract and introduction.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We have discussed the limitations and future work in Appendix E.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: We have provided the full set of assumptions and a complete (and correct) proof in the Method Section and Appendix A.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: We have shown our experiment results in the Experiment Section.
+We have submitted our source code in the Supplementary Material.
+We will upload our code and data to GitHub upon acceptance.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We have submitted our source code in the Supplementary Material. We will upload our code and data to GitHub upon acceptance.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We have shown our experimental settings and implementation details in Appendix D.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: Our reported results are all averaged over several random seeds. We have reported the standard deviation of the results in Table 2.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We have provided the compute resource in Appendix D.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The research conducted in the paper conforms, in every respect, with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We have discussed both potential positive societal impacts and negative societal impacts of the work in the Broader Impact Section.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper poses no such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: The datasets we used in the paper are all public datasets. Please refer to Appendix D for details of datasets.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: The paper does not release new assets.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v3

@@
 Our results provide fresh perspectives on designing spiking Transformers to advance their sequential modeling capability, thereby expanding their applicability across various domains.
+Our code is available at https://github.com/microsoft/SeqSNN.
 1 Introduction
@@
 As a result, the attention map $\mathbf{AttnMap}\in\mathbb{N}_{0}^{T\times L\times L}$, where $\mathbb{N}_{0}$ denotes the set of non-negative integers.
-The outputs of the SSA, as well as $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, are all spike matrices containing only values of 00 and $1$.
+The outputs of the SSA, as well as $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, are all spike matrices containing only values of 0 and $1$.
 The parameters $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, $\mathbf{W}_{V}$, and $\mathbf{W}$ are all learnable parameters.
@@
 Here, $\mathbf{R}_{i,j}$ represents the relative positional bias between the $i$-th query and the $j$-th key positions.
-A critical aspect of RPE is its adherence to distance consistency: the magnitude of $\mathbf{R}_{i,j}$ is determined exclusively by the relative positional offset $|i-j|$, ensuring that the model systematically differentiates between proximal and distant tokens.
+Beyond additive bias terms, another widely adopted form of RPE leverages relative positional embeddings directly in the attention computation, where query–position and key–position interactions are parameterized separately.
+For example, RoPE [14] can be expressed as
+$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\underbrace{\text{Softmax}\left(\frac{(\mathbf{Q}\mathbf{R}_{i})\cdot(\mathbf{K}\mathbf{R}_{j})^{\operatorname{T}}}{\sqrt{d_{k}}}\right)}_{\mathbf{AttnMap}}\cdot\mathbf{V},$ (6)
+where $\mathbf{R}_{i}$ and $\mathbf{R}_{j}$ are position-dependent rotation operators applied to the $i$-th query and $j$-th key vectors, respectively.
+A critical aspect of RPE is its adherence to distance consistency: the magnitude of $\mathbf{R}_{i,j}$ is determined exclusively by the relative positional offset |ij||i-j|, ensuring that the model systematically differentiates between proximal and distant tokens.
 This property enhances the model’s capacity to capture long-range dependencies and generalize across variations in sequence length and structure.
@@
 Formally, for two binary strings $A$ and $B$ of length $m$,
-$\small d_{H}(A,B)=\sum_{i=1}^{m}\delta(A_{i},B_{i}),\quad\text{where}\quad\delta(A_{i},B_{i})=\begin{cases}1&\text{if }A_{i}\neq B_{i},\\ 0&\text{otherwise}.\end{cases}$ (6)
+$d_{H}(A,B)=\sum_{i=1}^{m}\delta(A_{i},B_{i}),\quad\text{where}\quad\delta(A_{i},B_{i})=\begin{cases}1&\text{if }A_{i}\neq B_{i},\\ 0&\text{otherwise}.\end{cases}$ (7)
 Hamming distance is suitable for measuring the relative distances when $\mathbf{Q}$ and $\mathbf{K}$ are spike matrices.
@@
 For a non-negative integer $x$, the standard binary reflected Gray Code $G(x)$ is defined by the following bitwise operation:
-$\small G(x)=x\oplus\left(x\gg 1\right),$ (7)
+$G(x)=x\oplus\left(x\gg 1\right),$ (8)
 where $\oplus$ denotes the bitwise XOR operation, and $\gg$ denotes the arithmetic right shift.
@@
 4.3 Gray-PE
-We propose that the Gray Code can serve as an approximate approach to relative positional encoding for spiking Transformers. This is supported by the following Theorem 1:
+We propose that the Gray Code can serve as an approximate approach to relative positional encoding for spiking Transformers. This is supported by the following Theorem˜1:
 Theorem 1.
@@
 Specifically, $\forall$ position $i$, we have: $d_{H}(G(i),G(i+2^{n}))=\begin{cases}1&\text{if }n=0,\\ 2&\text{if }n\geq 1.\end{cases}$ (10)
-As illustrated in Figure 2 (b), the Hamming distance $d_{H}(0,1)$ and $d_{H}(1,2)$ both equal $1$ because their relative distance is $1$, i.e., $02^{n},n=$ 0.
+As illustrated in Figure 2 (b), the Hamming distance $d_{H}(0,1)$ and $d_{H}(1,2)$ both equal $1$ because their relative distance is $1$, i.e., $2^{n},n=0$.
 Similarly, $d_{H}(0,2)=d_{H}(1,3)$, and $d_{H}(0,4)=d_{H}(1,5)$, as their relative distances are the power of $2$.
@@
 where $G(\cdot)$ represents the function that converts integers into their binary Gray Code representations.
-The vector $\bm{l}$ denotes an array of position indexes, specifically $[0,1,2,\dots,L-1]$, where $L$ is the sequence length of $\mathbf{Q}$ and $\mathbf{K}$.
+The vector $\bm{l}$ denotes an array of position indexes, specifically $0,1,2,\dots,L-1]$, where $L$ is the sequence length of $\mathbf{Q}$ and $\mathbf{K}$.
 $\parallel$ denotes concatenation on the channel dimension $D$.
-Notably, the binary nature of Gray Code (comprising only 00 and $1$) aligns intrinsically with the spike-based computation paradigm, avoiding the need for floating-point operations that impose significant implementation overhead on neuromorphic hardware.
+Notably, the binary nature of Gray Code (comprising only 0 and $1$) aligns intrinsically with the spike-based computation paradigm, avoiding the need for floating-point operations that impose significant implementation overhead on neuromorphic hardware.
 4.4 Log-PE
@@
 Specifically, we simulate Equation 5 and follow ALiBi [13] to directly add a pre-assigned relative position map, denoted as $\mathbf{R}_{ij}$, to the attention map produced by SSA:
-$\mathbf{AttnMap}=\left(\sum_{i=0}^{D}\lnot(\mathbf{Q}\oplus\mathbf{K})\right)+\mathbf{R}_{i,j},\;\text{where}\;\;\mathbf{R}_{i,j}=\begin{bmatrix}R_{i,j}\end{bmatrix}=\begin{bmatrix}\lceil\log_{2}(\frac{L-1}{|i-j|+1})\rceil\end{bmatrix}.$ (11)
-Here, $\lceil$.$\rceil$ . ⌉ denotes the round-up function, $L$ is the sequence length, and $i,j$ is position indices.
+$\mathbf{AttnMap}=\left(\sum_{i=0}^{D}\lnot(\mathbf{Q}\oplus\mathbf{K})\right)+\mathbf{R}_{i,j},\;\text{where}\;\;\mathbf{R}_{i,j}=\begin{bmatrix}R_{i,j}\end{bmatrix}=\begin{bmatrix}\lceil\log_{2}(\frac{L-1}{|i-j|+1})\rceil\end{bmatrix}$.
+(12)
+Here, $\lceil.\rceil$ denotes the round-up function, $L$ is the sequence length, and $i,j$ is position indices.
 Figure 2 (c) shows an illustration of Log-PE.
@@
 Theoretically, if we set the $R_{i,j}$ as $\frac{L-1}{|i-j|+1}$, we could obtain a complete RPE for the spiking Transformers.
-However, we choose not to pursue this solution, because for long sequence lengths $L$, the large values of $\frac{L-1}{|i-j|+1}$ would catastrophically overshadow the original spiking attention activations (See Appendix B).
+However, we choose not to pursue this solution, because for long sequence lengths $L$, the large values of $\frac{L-1}{|i-j|+1}$ would catastrophically overshadow the original spiking attention activations (See Appendix˜B).
 Therefore, using the logarithmic form $R_{i,j}$ represents a compromise that balances the values between the spiking attention map and complete-RPE, while partially capturing relative position information.
@@
 The details of these datasets, metrics, and training hyperparameters are provided in Appendix D.
-Table 1: Experimental results of time-series forecasting on $4$ benchmarks with various prediction lengths $6,24,48,96$.
+Table 1: Experimental results of time-series forecasting on 44 benchmarks with various prediction lengths $6,24,48,96$.
 “PE” stands for positional encoding.
@@
 Results highlighted with shading are ours.
-All results are averaged across $3$ random seeds.
+All results are averaged across 33 random seeds.
 Models PE Metric Metr-la ($L=12$) Pems-bay ($L=12$) Solar ($L=168$) Electricity ($L=168$) Avg.
-Spike Type $6$ $24$ $48$ $96$ $6$ $24$ $48$ $96$ $6$ $24$ $48$ $96$ $6$ $24$ $48$ $96$
-Transformer w/ RoPE R R2$\uparrow$ .$\bf.729$$\bf.560$$\bf.416$$\bf.306$$\bf.787$$730$$\bf.694$$\bf.676$$951$$854$$\bf.763$$\bf.720$$\bf.984$$\bf.978$$\bf.974$$\bf.968$$\bf.756$
+Spike Type 66 $24$ $48$ $96$ 66 $24$ $48$ $96$ 66 $24$ $48$ $96$ 66 $24$ $48$ $96$
 […]
-Transformer w/ Sin-PE A R2$\uparrow$ .$727$$554$$413$$284$$785$$\bf.734$$688$$673$$\bf.953$$\bf.858$$759$$718$$978$$975$$972$$964$752.752.752.752
 […]
-Spikformer w/ Conv-PE (Original) A R2$\uparrow$ .$713$$527$$399$$267$$773$$697$$686$$667$$929$$828$$744$$674$$959$955.955.955.955 .955.955.955.955 .$954$733.733.733.733
 […]
-Spikformer w/ RoPE R2$\uparrow$ .$699$$493$$390$$243$$768$$699$$680$$664$$911$$820$$714$$644$$954$$951$$949$$940$720.720.720.720
 […]
-Spikformer w/ CPG-PE R2$\uparrow$ .$726$$526$$419$$287$$780$$712$$690$$666$$937$$833$$757$$707$$972$$970$$966$$960$744.744.744.744
 […]
-Spikformer-XNOR w/ Conv-PE A R2$\uparrow$ .$718$$531$$405$$269$$771$$693$$690$$665$$928$$829$$740$$669$$960$$957$$955$$953$733.733.733.733
 […]
-R2$\uparrow$ .$728$$\bf.544$$414$$\bf.295$$782$$\bf.724$$\bf.694$$\bf.673$$\bf.936$$840$$756$$710$$974$$972$$966$$962$748.748.748.748
 […]
-R2$\uparrow$ .$\bf.735$$535$$\bf.424$$290$$\bf.789$$717$$691$$670$$933$$\bf.841$$\bf.758$$\bf.734$$\bf.978$$\bf.974$$\bf.968$$\bf.964$$\bf.750$
 […]
-SDT-V1 w/ Conv-PE (Original) R2$\uparrow$ .$689$$517$$409$$253$$769$$700$$647$$630$$917$$819$$723$$655$$956$$952$$949$$950$721.721.721.721
 […]
-SDT-V1 w/ CPG-PE R2$\uparrow$ .$701$$525$$\bf.418$$257$$778$$\bf.716$$660$$\bf.656$$919$$\bf.820$$710$$644$$963$$960$$958$$952$727.727.727.727
 […]
-R2$\uparrow$ .$\bf.714$$\bf.531$$415$$\bf.265$$\bf.784$$709$$\bf.672$$654$$\bf.921$$\bf.820$$\bf.730$$\bf.674$$\bf.972$$\bf.968$$\bf.963$$\bf.957$$\bf.734$
 […]
-QKFormer w/ Conv-PE (Original) A R2$\uparrow$ .$717$$513$$376$$246$$767$$706$$681$$654$$920$$748$$512$$416$$970$$967$$963$$958$695.695.695.695
 […]
-QKFormer w/ CPG-PE A R2$\uparrow$ .$740$$554$$419$$276$$783$$714$$702$$660$$922$$754$$702$$604$$977$$969$$968$$963$732.732.732.732
 […]
-R2$\uparrow$ .$\bf.742$$\bf.551$$\bf.418$$\bf.274$$799$$\bf.715$$691$$\bf.674$$927$$817$$710$$691$$974$$970$$968$$965$742.742.742.742
 […]
-R2$\uparrow$ .$\bf.742$$541$$416$$265$$\bf.801$$710$$\bf.707$$661$$\bf.928$$\bf.818$$\bf.748$$\bf.698$$\bf.978$$\bf.974$$\bf.972$$\bf.966$$\bf.746$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 […]
@@
 We follow the SeqSNN [6] framework to conduct time-series forecasting experiments.
-Specifically, we take Spikformer [4], Spike-driven Transformer (SDT) V1 [5], and the current visual state-of-the-art (SOTA) model, QKFormer [20], as the backbone architectures.
+Specifically, we take Spikformer [4], Spikingformer [31], Spike-driven Transformer (SDT) V1 [5], and the current visual state-of-the-art (SOTA) model, QKFormer [20], as the backbone architectures.
 We modify the SSA mechanism as outlined in Section 4.2 to create two variants: Spikformer-XNOR and QKFormer-XNOR.
@@
 (1) Directly applying RPE methods to spiking Transformers is ineffective.
-Specifically, Spikformer with rotary position embedding (RoPE) exhibits poor performance across all benchmarks.
+Specifically, Spikformers that are directly equipped with RoPE or ALiBi exhibit poor performance across all benchmarks.
 As discussed in Section 1, we argue that this limitation stems from the binary nature of spiking neurons during the computation of $\mathbf{Q}$ and $\mathbf{K}$, which makes it difficult to disentangle positional information from sparse spiking activations.
@@
 Our experimental setup strictly adheres to the methodology outlined in [7], and the results are shown in Table 2.
-Table 2: Accuracy (%) on $6$ text classification benchmarks. Note that QKFormers fail to converge in the text classification task. Experimental results are averaged across $5$ random seeds.
+Table 2: Accuracy (%) on 66 text classification benchmarks. Note that QKFormers fail to converge in the text classification task. Experimental results are averaged across 55 random seeds.
 Model PE Param(M) English Dataset (Length = $128$) Chinese Dataset (Length = $32$) Avg.
 Spike Type MR SST-2 Subj SST-5 ChnSenti Waimai
-Fine-tuned BERT A $109.8$ $\bm{87.63}$±$\pm 0.18$ $\bm{92.31}$±$\pm 0.17$ $\bm{95.90}$±$\pm 0.16$ $\bm{50.41}$±$\pm 0.13$ $\bm{89.48}$±$\pm 0.16$ $\bm{90.27}$±$\pm 0.13$ $\bm{84.33}$
-Spikformer w/o PE $109.8$ $75.87$±$\pm 0.35$ $81.71$±$\pm 0.31$ $91.60$±$\pm 0.30$ $41.84$±$\pm 0.39$ $85.62$±$\pm 0.25$ $86.87$±$\pm 0.28$ $77.25$
-Spikformer w/ CPG-PE A $110.4$ $82.42$±$\pm 0.42$ $82.90$±$\pm 0.33$ $92.50$±$\pm 0.25$ $43.62$±$\pm 0.36$ $86.54$±$\pm 0.26$ $\bf 88.49$±$\pm 0.29$ $79.41$
-Spikformer-XNOR w/o PE $109.8$ $75.80$±$\pm 0.40$ $81.74$±$\pm 0.40$ $91.50$±$\pm 0.29$ $41.88$±$\pm 0.38$ $85.64$±$\pm 0.31$ $86.66$±$\pm 0.33$ $77.20$
-Spikformer-XNOR w/ Gray-PE R $109.8$ $83.73$±$\pm 0.45$ $84.52$±$\pm 0.39$ $92.50$±$\pm 0.33$ $44.06$±$\pm 0.48$ $87.41$±$\pm 0.36$ $88.40$±$\pm 0.30$ $80.11$
-Spikformer-XNOR w/ Log-PE R $109.8$ $\bf 83.88$±$\pm 0.40$ $\bf 84.64$±$\pm 0.37$ $\bf 92.80$±$\pm 0.30$ $\bf 44.52$±$\pm 0.43$ $\bf 87.95$±$\pm 0.34$ $88.46$±$\pm 0.28$ $\bf 80.38$
+Fine-tuned BERT A $109.8$ $\bm{87.63}$$\pm 0.18$ $\bm{92.31}$$\pm 0.17$ $\bm{95.90}$$\pm 0.16$ $\bm{50.41}$$\pm 0.13$ $\bm{89.48}$$\pm 0.16$ $\bm{90.27}$$\pm 0.13$ $\bm{84.33}$
+[…]
+[…]
+[…]
+[…]
+[…]
 Based on the results in Table 2, it is evident that our proposed Gray-PE and Log-PE significantly outperform the other spiking positional encoding methods across several key benchmarks.
@@
 The results with shading are ours.
-Results are averaged across $4$ random seeds.
+Results are averaged across 44 random seeds.
 Model PE CIFAR10 CIFAR10-DVS CIFAR100 Tiny-ImageNet Avg.
@@
 Specifically, Gray-PE 2D achieves an average accuracy improvement of $0.44\%$ over Gray-PE 1D.
-Furthermore, we present the image classification performance of the state-of-the-art QKFormer integrated with our proposed RPE methods in Appendix C.
+Furthermore, we present the image classification performance of the state-of-the-art QKFormer integrated with our proposed RPE methods in Appendix˜C.
 5.5 Capability of Processing Long Sequences
@@
 We train the Spikformer model using various positional encoding strategies on these datasets, and present the results in Table 4.
-Table 4: Accuracy (%) on $2$ long text classification benchmarks. We set the sentence length to $1024$ for AGNEWS and $2048$ for IMDB.
+Table 4: Accuracy (%) on 22 long text classification benchmarks. We set the sentence length to $1024$ for AGNEWS and $2048$ for IMDB.
 Model PE AGNEWS IMDB Avg.
@@
 This outcome is expected, as Log-PE is specifically designed to accommodate long-range dependencies.
+5.6 Discussion on Hardware-Friendliness and Computing Efficiency
+Table 5: Evaluation of both time consumption and GPU memory usage for SNNs on Electricity dataset.
+Model Time Consumption GPU Memory Usage
+s/epoch MB
+Spikformer (Original) $137.48$ $10572.56$
+Spikformer-XNOR $139.66$ $10608.32$
+Spikformer w/ CPG-PE $140.56$ $10963.88$
+Although traditional SSA benefits from highly optimized matrix multiplication (GEMM) on GPUs, we would like to clarify that our XNOR-based SSA also retains computational efficiency for the following reasons: First, the core of XNOR-based SSA relies on XNOR and bit-count operations, which are natively supported by digital hardware and neuromorphic processors.
+These are much cheaper than floating-point multiplications and additions in terms of energy and hardware complexity.
+Secondly, many neuromorphic accelerators (e.g., Loihi [35], TrueNorth [36]) natively support spike-based bitwise logic, making our XNOR mechanism better aligned with the target deployment platform than conventional floating-point matrix products.
+Lastly, while matrix multiplication benefits from BLAS acceleration, XNOR and summation over dimensions are also highly parallelizable, and can be efficiently implemented using tensor intrinsics (e.g., bitwise_xnor, popcount, reduce_sum).
+We benchmarked both time consumption and GPU memory usage for SNNs in a time-series forecasting task, mainly on the Electricity dataset with 24 of horizon length, as shown in 5.
+For more analysis on the hardware-friendliness of Log-PE, please refer to the Appendix˜E.
 5.7 Analysis and Ablation
@@
 Consider that: If the number of bits used for encoding relative positions in Gray Code is $b$, then the total number of unique encodings possible is $2^{b}$.
-We set the maximum sequence length is $L$, so relative distances range from 00 to $L-1$.
+We set the maximum sequence length is $L$, so relative distances range from 0 to L1L-1.
 According to the pigeonhole principle, if $L-1>2^{b}$, there will be at least two distances that are represented identically.
-This issue can be mitigated by increasing $b$ to cover the range of relative distances up to $L-1$.
-From Figure 3 (a), we observe that for long-sequence datasets, such as Solar and Electricity (Length $=168$), the number of bits should be at least $7$ to avoid Gray-PE missing relative positional information.
-However, for shorter datasets like Metr-la (Length $=12$) and ChnSenti (Length $=32$), $5$ bits are sufficient.
+This issue can be mitigated by increasing $b$ to cover the range of relative distances up to L1L-1.
+From Figure 3 (a), we observe that for long-sequence datasets, such as Solar and Electricity (Length =168=168), the number of bits should be at least $7$ to avoid Gray-PE missing relative positional information.
+However, for shorter datasets like Metr-la (Length =12=12) and ChnSenti (Length =32=32), 55 bits are sufficient.
 6 Conclusion
@@
 These empirical results, together with theoretical analysis of the proposed RPE methods, highlight the potential to enhance the versatility and applicability of SNNs across various domains.
-Future work and limitations are discussed in Appendix E.
+Future work and limitations are discussed in Appendix F.
+Broader Impact
+This work aims to advance the field of spiking neural networks.
+We hope our work can open new avenues for embedding relative positional encoding within SNNs, thereby expanding their applicability across a wide range of domains.
+We do not see negative societal impacts of this work.
+Acknowledge
+The authors would like to thank the anonymous reviewers for their valuable comments. This work was partially supported by the National Natural Science Foundation of China (No. 62076068).
 Appendix A Proof of Theorem 1
-This section provides a detailed mathematical proof of Theorem 1. We use the standard Reflected Binary Code (RBC) $G(x)$ as our Gray Code:
+This section provides a detailed mathematical proof of Theorem˜1. We use the standard Reflected Binary Code (RBC) $G(x)$ as our Gray Code:
 Definition 1.
@@
 $G(a)\oplus G(b)=2^{n}\oplus 2^{n-1}.$ (20)
-This value has exactly two bits set (at positions $n$ and $n-1$), so the Hamming distance is 2.
+This value has exactly two bits set (at positions $n$ and n1n-1), so the Hamming distance is 2.
 Subcase B: Bit $n$ in $a$ is 1
-Then adding $2^{n}$ to $a$ causes a carry from bit $n$ upwards. Let $c$ be the smallest index greater than $n$ such that bit $c$ in $a$ is 00; bits $n$ through $c-1$ are all $1$. Then:
+Then adding $2^{n}$ to $a$ causes a carry from bit $n$ upwards. Let $c$ be the smallest index greater than $n$ such that bit $c$ in $a$ is 0; bits $n$ through c1c-1 are all $1$. Then:
 $\Delta_{1}=a\oplus(a+2^{n})=\sum_{i=n}^{c}2^{i}=\underbrace{1\dots 1}_{c-n+1}\underbrace{0\dots 0}_{n}\ {}_{(2)}$, (21)
@@
 $\Delta_{2}=(a\gg 1)\oplus((a+2^{n})\gg 1)=\sum_{i=n-1}^{c-1}2^{i}=\underbrace{1\dots 1}_{c-n+1}\underbrace{0\dots 0}_{n-1}\ {}_{(2)}$, (22)
-has ones at bits $n-1$ through $c-1$. Thus,
+has ones at bits n1n-1 through c1c-1. Thus,
 $G(a)\oplus G(b)=\Delta_{1}\oplus\Delta_{2}=(\sum_{i=n}^{c}2^{i})\oplus(\sum_{i=n-1}^{c-1}2^{i})=2^{c}+2^{n}$ (23)
-has ones at only positions $c$ and $n-1$, and all other bits are canceled due to alignment. The result has exactly two bits set, so the Hamming distance is $2$.
+has ones at only positions $c$ and n1n-1, and all other bits are canceled due to alignment. The result has exactly two bits set, so the Hamming distance is $2$.
 Combining all cases, we conclude that for any non-negative integer $n$:
@@
 In this section, we compare the performance of vanilla spiking Transformers with Log-PE, XNOR variants with Log-PE, and models equipped with complete relative positional encoding (C-RPE).
-As discussed in Section 4.4, C-RPE is implemented by setting $R_{i,j}=\frac{L-1}{|i-j|+1}$ and adding it directly to the attention scores.
+As discussed in Section˜4.4, C-RPE is implemented by setting $R_{i,j}=\frac{L-1}{|i-j|+1}$ and adding it directly to the attention scores.
 Table S1: Ablation study on XNOR and Log-PE.
@@
 denotes failure to converge.
-Model (Prediction Length $=24$) Metr-la ($L=12$) Electricity ($L=168$)
+Model (Prediction Length =24=24) Metr-la ($L=12$) Electricity ($L=168$)
 R2 $\uparrow$ RSE $\downarrow$ R2 $\uparrow$ RSE $\downarrow$
-Spikformer w/ Log-PE .$484$$763$$710^{*}.$ $1.03^{*}$
-Spikformer-XNOR w/ Log-PE .$\bf.535$$\bf.719$$\bf.974$$\bf.300$
-Spikformer-XNOR w/ C-RPE .$158^{*}.$$967^{*}.$$710^{*}.$ $1.03^{*}$
-QKFormer w/ Log-PE .$475$$788$$710^{*}.$ $1.03^{*}$
-QKFormer-XNOR w/ Log-PE .$\bf.541$$\bf.715$$\bf.974$$\bf.285$
-QKFormer-XNOR w/ C-RPE .$430$$824$$710^{*}.$ $1.03^{*}$
-As shown in Table S1, both Spikformer and QKFormer with C-RPE perform significantly worse than their Log-PE counterparts, with some variants failing to converge entirely.
+Spikformer w/ Log-PE .484.484 .763.763 $710^{*}$ $1.03^{*}$
+Spikformer-XNOR w/ Log-PE $\bf.535$ $\bf.719$ $\bf.974$ $\bf.300$
+Spikformer-XNOR w/ C-RPE $158^{*} .$$967^{*} .$$710^{*}$ $1.03^{*}$
+QKFormer w/ Log-PE .475.475 .788.788 $710^{*}$ $1.03^{*}$
+QKFormer-XNOR w/ Log-PE $\bf.541$ $\bf.715$ $\bf.974$ $\bf.285$
+QKFormer-XNOR w/ C-RPE .430.430 .824.824 $710^{*}$ $1.03^{*}$
+As shown in Table˜S1, both Spikformer and QKFormer with C-RPE perform significantly worse than their Log-PE counterparts, with some variants failing to converge entirely.
 This degradation is attributed to the overly large positional encodings disrupting the training dynamics.
@@
 “R” denotes relative PE, while “A” denotes absolute PE.
-Results are averaged across $4$ random seeds.
+Results are averaged across 44 random seeds.
 Model PE CIFAR10 CIFAR10-DVS CIFAR100 Tiny-ImageNet Avg.
@@
 $R^{2}$ $=\frac{1}{MCL}\sum_{m=1}^{M}\sum_{c=1}^{C}\sum_{l=1}^{L}\left[1-\frac{(Y^{m}_{c,l}-\hat{Y}^{m}_{c,l})^{2}}{(Y^{m}_{c,l}-\bar{Y}_{c,l})^{2}}\right],$ (25)
-$\mathrm{RSE}$ $=\sqrt{\frac{\sum_{m=1}^{M}||\mathbf{Y}^{m}-\hat{\mathbf{Y}}^{m}||^{2}}{\sum_{m=1}^{M}||\mathbf{Y}^{m}-\bar{\mathbf{Y}}||^{2}}}.$ (25)
+[…]
 In these formulas, $M$ represents the size of the test set, $C$ denotes the number of channels, and $L$ signifies the length of the predictions.
@@
 We construct all Spikformer as $2$ blocks, setting the feature dimension as $256$, and the hidden feature dimension in FFN as $1024$.
-As for QKFormer, we set the block number as $4$, $2$ of which are QK blocks and the other $2$ are Spikformer blocks.
+As for QKFormer, we set the block number as $4$, 22 of which are QK blocks and the other $2$ are Spikformer blocks.
 Training Hyper-parameters
@@
 As the training epochs are quite big ($300$ or $400$ epochs) in their settings, we choose to use one 80G-A100 GPU, and it takes about $3$ hours to conduct a single experiment, on average.
-Appendix E Limitations and Future Work
-E.1 Limitations
+Appendix E Analysis on Hardware-Friendliness of Log-PE
+First, Log-PE doesn’t need to perform logarithmic operations directly on hardware during inference.
+Specifically, the relative position bias is defined as $\mathbf{R}_{i,j}=\begin{bmatrix}\lceil\log_{2}\left(\frac{L-1}{|i-j|+1}\right)\rceil\end{bmatrix}$, where $\lceil\cdot\rceil$ denotes the ceiling (round-up) function, and $L$ is the maximum sequence length.
+Since this bias depends only on the relative positions and the predefined sequence length, the entire bias matrix can be computed offline and stored ahead of time, eliminating the need for any runtime computation.
+Secondly, even if one wishes to compute the logarithmic transformation on hardware, this can be efficiently achieved using a lookup table (LUT) implementation.
+Given an unsigned integer input of $N$ bits, we partition the input range into $K$ intervals.
+Each interval is approximated using a piecewise linear function $y=ax+b$, with the parameters $(a,b)$ stored in the LUT.
+The total LUT storage cost is: $K\cdot(N+2P)\ \text{bits}\approx\frac{K\cdot(N+2P)}{8}\ \text{bytes},$ where $P$ is the bit width of the parameters.
+This strategy is similar to existing SNN approximations for exponential/leaky functions and has been successfully deployed in many types of neuromorphic chips, such as Intel Loihi [35].
+Hence, the hardware implementation of Log-PE is efficient, low-cost, and practically feasible.
+Appendix F Limitations and Future Work
+F.1 Limitations
 Despite the promising enhancements introduced by our relative positional encoding method for spiking Transformers, several limitations must be acknowledged.
@@
 Furthermore, our evaluations have been confined to specific applications such as time series forecasting, text classification, and image patch classification, which may not fully capture the method’s performance in other domains, such as object detection [44] and real-world geometry representation [45].
-E.2 Future Work
+F.2 Future Work
 Future work should focus on optimizing the Gray Code-based RPE to enhance its scalability and efficiency, enabling its deployment in larger and more intricate SNN models.
@@
   8. Answer: [Yes]
-  9. Justification: We have discussed the limitations and future work in Appendix E.
+  9. Justification: We have discussed the limitations and future work in Appendix F.
   10. Guidelines:
@@
   33. Answer: [Yes]
-  34. Justification: Our reported results are all averaged over several random seeds. We have reported the standard deviation of the results in Table 2.
+  34. Justification: Our reported results are all averaged over several random seeds. We have reported the standard deviation of the results in Table˜2.
   35. Guidelines:
@@
   58. Answer: [Yes]
-  59. Justification: The datasets we used in the paper are all public datasets. Please refer to Appendix D for details of datasets.
+  59. Justification: The datasets we used in the paper are all public datasets. Please refer to Appendix˜D for details of datasets.
   60. Guidelines: